### v2

# Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

Abstract
Crossed random effects structures arise in many scientific contexts. They raise severe computational problems with likelihood computations scaling like $N^{3/2}$ or worse for $N$ data points. In this paper we develop a new composite likelihood approach for crossed random effects probit models. For data arranged in $R$ rows and $C$ columns, the likelihood function includes a very difficult $R+C$ dimensional integral. The composite likelihood we develop uses the marginal distribution of the response along with two hierarchical models. The cost is reduced to $\mathcal{O}(N)$ and it can be computed with $R+C$ one dimensional integrals. We find that the commonly used Laplace approximation has a cost that grows superlinearly. We get consistent estimates of the probit slope and variance components from our composite likelihood algorithm. We also show how to estimate the covariance of the estimated regression coefficients. The algorithm scales readily to a data set of five million observations from Stitch Fix with $R+C>700{,}000$.
Keywords: Composite Likelihood; High-Dimensional Binary Data; Probit Regression; Scalability.
1 Introduction
In this paper we develop a scalable approach to mixed effects modeling with a probit link and a crossed random effects error structure. The common notation for mixed effects models incorporates the fixed and random effects through a formula such as $X\bm{\beta}+Z\bm{b}$ using matrices $X$ and $Z$ of known predictors and unknown coefficient $\bm{\beta}$ and $\bm{b}$ where $\bm{b}$ happens to be random. This formulation is simple and elegant but it hides an important practical difference. When the random effects have a hierarchical structure then computations usually scale well to large problems. However, when the random effects have a crossed structure, then computations can easily grow superlinearly, which greatly restricts the use of these models. For instance, computing the maximum likelihood estimator in a Gaussian model with crossed random effects by the usual algorithms includes linear algebra with a cost that grows at least as fast as $N^{3/2}$ given $N$ data points (Gao and Owen, 2020), while remaining $\mathcal{O}(N)$ in the hierarchical setting. The binary regression setting we consider here has the further difficulty of a very high dimensional integral that is not present for a model with a Gaussian response. Maximizing the likelihood with the expectation-maximization algorithm raises the same intractability problems as direct likelihood maximization because the integral appearing in the expectation step has the same very high dimension as the integral appearing in the likelihood.
In an era where the size $N$ of data sets is growing rapidly it is not possible to use algorithms with a cost of $\Omega(N^{3/2})$. Ideally the cost should be $\mathcal{O}(N)$. We are motivated by electronic commerce problems with large data sets. A company might have customers $i=1,\dots,R$, to which it sells items $j=1,\dots,C$. It would then be interested in modeling how a response $Y_{ij}$ depends on some predictors $\bm{x}_{ij}\in\mathbb{R}^{p}$. If they do not account for the fact that $Y_{ij}$ and $Y_{is}$ are correlated due to a common customer $i$ or that $Y_{ij}$ and $Y_{rj}$ are correlated due to a common item $j$, they will not get efficient estimates of the coefficients in their model. Much more seriously, they will not get reliable standard errors for those estimates. In a setting where accounting for random effects is computationally impossible we expect that many users will simply ignore them, getting very naive variance estimates and finding too many things significant.
A typical feature of data in our motivating applications is a very sparse sampling. Only $N\ll RC$ of the possible $(\bm{x}_{ij},Y_{ij})$ values are observed. There is generally no simple structure in the pattern of which $(i,j)$ pairs are observed. While we use electronic commerce as our example, large sparsely sampled crossed data sets arise in other areas too.
One of the main open questions in generalized linear mixed models, like the ones we consider here, is at what rate do the estimators of the model parameters converge? Intuitively one might expect $\mathcal{O}(N^{-1/2})$ or $\mathcal{O}(\min\{R,C\}^{-1/2})$. In some settings estimators for different parameters converge at different rates. See Jiang (2013) for some discussion. It was only comparatively recently that consistency of maximum likelihood estimation was proved in some generalized linear mixed models. This is also due to Jiang (2013).
In this paper we focus on binary responses, $Y_{ij}\in\{0,1\}$. In commerce applications, such responses may be more important than real-valued ones. Examples include whether customer $i$ purchased item $j$, or returned it or reported satisfaction with it. In addition to the numerical algebra that underlies the $\Omega(N^{3/2})$ cost of standard algorithms this setting brings a second difficulty of an $R+C$ dimensional integral over random effect values in the likelihood. It is very common to consider a logistic regression for such binary responses. In this paper we consider a probit regression. The probit model is especially suitable because it is typical to use Gaussian models for the random effects. The probit and logit link functions are nearly proportional outside of tail regions (Agresti, 2002, pp. 246–247), and so they often give similar results. We have some further remarks about the logistic case in our conclusions.
We consider two crossed random effects. There is a vector $\bm{a}\in\mathbb{R}^{R}$ with elements $a_{i}$ and another vector $\bm{b}\in\mathbb{R}^{C}$ with elements $b_{j}$. Conditionally on $\bm{a}$ and $\bm{b}$, the $Y_{ij}$ are independent with
$\Pr(Y_{ij}=1\!\mid\!\bm{a},\bm{b})=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j}),$ (1)
where $\Phi(\cdot)$ is the standard normal cumulative distribution function. In these models, random effects are typically assumed to be uncorrelated, $\bm{a}\sim\mathcal{N}(0,\sigma^{2}_{A}I_{R})$ independently of $\bm{b}\sim\mathcal{N}(0,\sigma^{2}_{B}I_{C}),$ where $I_{n}$ is the $n\times n$ identity matrix, see, for example, McCullagh and Nelder (1989, page 444). The probit model has a representation in terms of latent variables $\varepsilon_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,1)$, via
$Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j}+\varepsilon_{ij}>0\},$ (2)
where $1\{E\}$ is the indicator function of the event $E$. The probability (1) and the likelihoods derived from it are all conditional on the values of $\bm{x}_{ij}$. By combining (1) and (2) we find that marginally
$\Pr(Y_{ij}=1)=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})$ (3)
for $\bm{\gamma}=\bm{\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}$. The marginal model (3) can be fit in $\mathcal{O}(N)$ computation, but we then need estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to get the scale right. While $\mathrm{sign}(\bm{\gamma}_{k})=\mathrm{sign}(\beta_{k})$ for $k=1,\dots,p$ confidence intervals for $\bm{\gamma}_{k}$ based on model (3) would be naive if they did not account for the dependence among the responses.
We can similarly find that
$\Pr(Y_{ij}=1\!\mid\!\bm{a})$ $=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{A}+u_{i}),\quad\text{and}$ (4)
$\Pr(Y_{ij}=1\!\mid\!\bm{b})$ $=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{B}+v_{j})$ (5)
where $\bm{u}\sim\mathcal{N}(0,\tau^{2}_{A}I_{R})$ and $\bm{v}\sim\mathcal{N}(0,\tau^{2}_{B}I_{C})$ for $\tau^{2}_{A}=\sigma^{2}_{A}/(1+\sigma^{2}_{B})$, $\tau^{2}_{B}=\sigma^{2}_{B}/(1+\sigma^{2}_{A})$, $\bm{\gamma}_{A}=\bm{\gamma}(1+\tau^{2}_{A})^{1/2}$ and $\bm{\gamma}_{B}=\bm{\gamma}(1+\tau^{2}_{B})^{1/2}$. We fit models (3), (4) and (5) together to estimate our probit model, using the latter two to estimate $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Fitting models (4) and (5) involve simpler integrals than (1) since the latent variable representations of these models
$Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{A}+u_{i}+\varepsilon_{ij}>0\}\quad\text{and}\quad Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{B}+v_{j}+\varepsilon_{ij}>0\}$
have hierarchical (not crossed) error structures. This is where we are able to replace an $R+C$-dimensional integral by $R+C$ univariate ones.
We call our method ‘all-row-column’ (ARC). The name comes from model (3) that uses all the data at once, model (4) that combines likelihood contribution from within each row and model (5) that combines likelihood contributions from within each column. This approach is a new form of composite likelihood (Lindsay, 1988; Varin et al., 2011). An earlier version of composite likelihood that was applied to crossed random effects is discussed in Bellio and Varin (2005). They considered a standard pairwise likelihood which is however not scalable and thus inappropriate for our problem. Instead, our ARC method does not only consistently estimates $\bm{\beta}$, $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ but it is also scalable having $\mathcal{O}(N)$ cost.
An outline of this paper is as follows: Section 2 develops further notation and properties for the models we combine. Section 3 presents our ARC method and develops a robust sandwich estimate of the variance of the regression parameter estimates. Section 4 presents some simulations of the ARC method. We also compare it to an infeasible oracle method that estimates $\bm{\gamma}$ from (3) and then adjusts the estimate using the true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. A second comparison is to the most popular frequentist method for fitting generalized linear models with random effects, that is the first-order Laplace approximate maximum likelihood estimator described in Skaug (2002) and implemented in the popular R (R Core Team, 2024) package lme4 (Bates et al., 2015). Section 5 illustrates our proposed method on real electronic commerce data provided by Stitch Fix. Section 6 gives some bibliographic references focussing on the challenges that the crossed random effects setting brings. Section 7 has our conclusions. Finally, there is a supplement with proofs of consistent estimation, more references, consideration of pairwise composite likelihood and a more detailed discussion of the simulation results.
2 Probit Regression with Crossed Random Effects
As described in the introduction we have responses $Y_{ij}\in\{0,1\}$, covariates $\bm{x}_{ij}\in\mathbb{R}^{p}$ and random effects $\bm{a}\sim\mathcal{N}(0,\sigma^{2}_{A}I_{R})$, $\bm{b}\sim\mathcal{N}(0,\sigma^{2}_{B}I_{C})$. The random effects are independent of each other, and we condition on the observed values of the covariates $\bm{x}_{ij}$, so we then treat them as non-random. We write $\mathcal{S}\subset\{1,\dots,R\}\times\{1,\dots,C\}$ for the set of $(i,j)$ pairs where $(\bm{x}_{ij},Y_{ij})$ was observed. We also work conditionally on $\mathcal{S}$. In our motivating applications the pattern of observation/missingness could well be informative. Addressing that issue would necessarily require information from outside the data. Furthermore, the scaling problem is still the subject of current research in the non-informative missingness setting. Therefore we consider estimation strategies without taking account of missingness.
We choose to study the probit model (1) where, conditionally on $\bm{a}$ and $\bm{b}$, $Y_{ij}\in\{0,1\}$ are independent with $\Pr(Y_{ij}=1\!\mid\!\bm{a},\bm{b})=\Phi(\bm{x}_{ij}^{\mathsf{T}}\beta+a_{i}+b_{j})$. The full likelihood for $\bm{\theta}=(\bm{\beta}^{\mathsf{T}},\sigma^{2}_{A},\sigma^{2}_{B})^{\mathsf{T}}$ is the integral
$L(\bm{\theta})=\sigma_{A}^{-R}\sigma_{B}^{-C}\int_{\mathbb{R}^{R+C}}L(\bm{\beta}\!\mid\!\bm{a},\bm{b})\prod_{i=1}^{R}\varphi\Bigl{(}\frac{a_{i}}{\sigma_{A}}\Bigr{)}\prod_{j=1}^{C}\varphi\Bigl{(}\frac{b_{j}}{\sigma_{B}}\Bigr{)}\,\mathrm{d}\bm{a}\,\mathrm{d}\bm{b},$ (6)
where $\varphi(\cdot)$ is the standard normal probability density function, and $L(\bm{\beta}\!\mid\!\bm{a},\bm{b})$ is the conditional likelihood of $\bm{\beta}$ given the random effects. The conditional likelihood we need is
$L(\bm{\beta}\!\mid\!\bm{a},\bm{b})=\prod_{(i,j)\in\mathcal{S}}\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}-a_{i}-b_{j})^{1-y_{ij}}.$
An option to alleviate the computational cost of likelihood computations is provided by composite likelihood methods (Lindsay, 1988). Composite likelihoods are inference functions constructed by multiplying component likelihoods. See Varin et al. (2011) for a review. Several articles discussed composite likelihood estimation of models for dependent binary data; see, for example, Heagerty and Lele (1998), Cox and Reid (2004) or Kuk (2007). All these works consider a particular type of composite likelihood called pairwise likelihood that is built taking the product of bivariate marginal probabilities. Bellio and Varin (2005) investigated pairwise likelihood inference in logistic regression with crossed random effects. Their pairwise likelihood involves all the pairs of correlated observations, that is those pairs that share the row- or the column-random effect.
Pairwise likelihood is not attractive in this setting. The number of pairs of data points far exceeds the $\mathcal{O}(N)$ constraint. In a setting with $R$ rows, the average row has $N/R$ elements in it. If all the rows had that many elements then the number of pairs would be $\Omega(R(N/R)^{2})=\Omega(N^{2}/R)$. Unequal numbers of observations per row can only increase this count. As a result the cost must be $\Omega(\max(N^{2}/R,N^{2}/C))$. In our motivating settings, both $R$ and $C$ are $o(N)$ so the cost cannot then be $\mathcal{O}(N)$. A possible remedy to the lack of scalability would be to use only $\mathcal{O}(N)$ randomly sampled pairs of correlated observations. Although sampling might be a valid option, we decided to avoid this approach. First, we prefer estimates that do not depend on unnecessary randomness in the algorithm. Second, some preliminary simulations reported in Section S2 of the Supplementary Material indicate that the composite likelihood discussed in the next section satisfy the $\mathcal{O}(N)$ constraint with a statistical efficiency comparable to that of the pairwise likelihood without sampling of the pairs, which is however not feasible to compute in the high-dimensional applications that motivate this work. Third, we view our ARC method as new form of composite likelihood using all pairs, all triples and indeed all subsets of observations within each of the rows, and also within each of the columns. It thus captures all the data dependencies already included in pairwise likelihood and more while only requiring three estimating equations and $\mathcal{O}(N)$ cost.
3 The All-Row-Column Method
3.1 Three misspecified probit models
Our ARC approach for scalable inference in high-dimensional probit models with crossed-random effects combines estimates for three misspecified probit models. Each of them is constructed through omission of some random effects.
Consider the reparameterization $\bm{\psi}=(\bm{\gamma}^{\mathsf{T}},\tau^{2}_{A},\tau^{2}_{B})^{\mathsf{T}},$ whose components were already anticipated in Section 1, namely,
$\bm{\gamma}=\frac{\bm{\beta}}{(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}},\quad\tau^{2}_{A}=\frac{\sigma^{2}_{A}}{1+\sigma^{2}_{B}}\quad\text{and}\quad\tau^{2}_{B}=\frac{\sigma^{2}_{B}}{1+\sigma^{2}_{A}}.$ (7)
The proposed method begins with estimation of $\bm{\gamma}$ from the naïve model (3) that omits both of the random effects through maximization of the likelihood of the probit model
$L_{\text{all}}(\bm{\gamma})=\prod_{(i,j)\in\mathcal{S}}\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})^{1-y_{ij}}.$ (8)
Given $\hat{\bm{\gamma}}$, the maximizer of (8), we proceed with estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the two probit models (4) and (5) that each omit one of the random effects. Accordingly, parameter $\tau^{2}_{A}$ is estimated from the row-wise likelihood
$L_{\text{row}}(\tau^{2}_{A})=\tau_{A}^{-R}\prod_{i=1}^{R}\int_{\mathbb{R}}L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})\varphi\Bigl{(}\frac{u_{i}}{\tau_{A}}\Bigr{)}\,\mathrm{d}u_{i},$ (9)
where $L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})$ is the conditional likelihood of $\hat{\bm{\gamma}}_{A}=\hat{\bm{\gamma}}(1+\tau^{2}_{A})^{1/2}$ given $u_{i},$
$L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})=\prod_{j|i}\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{A}+u_{i})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{A}-u_{i})^{1-y_{ij}},$ (10)
where $j|i={\{j:(i,j)\in\mathcal{S}\}}$ is the set of indices $j$ such that $(\bm{x}_{ij},Y_{ij})$ is observed. Notice that the row-wise likelihood is a function of $\tau^{2}_{A}$ only, because we fix $\bm{\gamma}$ at the estimate $\hat{\bm{\gamma}}$ obtained from the maximization of the all likelihood at the previous step. Rows with a single observation do not contribute to estimation of $\tau^{2}_{A}$ because $\Pr(Y_{ij}=1)=\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}),$ which does not depend on $\tau^{2}_{A}.$
Computation of the row-wise likelihood requires one to approximate $R$ univariate integrals of a standard hierarchical probit model. Since $R$ is large in the applications that motivate this work, then accurate approximation of the one-dimensional integrals is crucial. Otherwise, the accumulation of approximation errors could induce serious biases in the estimation of $\tau^{2}_{A}$. Our intensive numerical studies indicated that accurate approximation of the integrals is obtained by the well-established adaptive Gaussian quadrature (Liu and Pierce, 1994) with a suitable choice of the number of quadrature nodes. Following Jin and Andersson (2020) and Stringer and Bilodeau (2022), we set the number of quadrature nodes equal to $k(R)=\lceil(3/2)\log_{2}(R)-2\rceil,$ a number rather larger than is typically considered in the GLMM literature. For example, in the Stitch Fix application discussed in Section 5, we have $R=741,221$ and therefore we have used $k=28$ quadrature nodes.
Reversing rows and columns, $\tau^{2}_{B}$ is estimated from the column-wise likelihood
$L_{\text{col}}(\tau^{2}_{B})=\tau_{B}^{-C}\prod_{j=1}^{C}\int_{\mathbb{R}}L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})\varphi\Bigl{(}\frac{v_{j}}{\tau_{B}}\Bigr{)}\,\mathrm{d}v_{j},$ (11)
where $L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})$ is the conditional likelihood of $\hat{\bm{\gamma}}_{B}=\hat{\bm{\gamma}}(1+\tau^{2}_{B})^{1/2}$ given $v_{j},$
$L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})=\prod_{i|j}\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{B}+v_{j})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{B}-v_{j})^{1-y_{ij}}$
with $i|j={\{i:(i,j)\in\mathcal{S}\}}.$ Finally, the estimates under the original parameterization are obtained by inverting the equations in (7),
$\hat{\bm{\beta}}=\hat{\bm{\gamma}}(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2},\quad\hat{\sigma}_{A}^{2}=\frac{\hat{\tau}_{A}^{2}(1+\hat{\tau}^{2}_{B})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}\quad\text{and}\quad\hat{\sigma}_{B}^{2}=\frac{\hat{\tau}_{B}^{2}(1+\hat{\tau}^{2}_{A})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}.$
Note that by the definitions in (7), $\tau^{2}_{A}\tau^{2}_{B}<1$. In our computations, we never encountered a setting where $\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}>1$ [should be $\geq$ not $>$] and so our estimates $\hat{\sigma}^{2}_{A}$ and $\hat{\sigma}^{2}_{B}$ were never negative. In case of a violation, we suggest taking $\hat{\sigma}^{2}_{A}=\hat{\sigma}^{2}_{B}=0$ as is commonly done for negative variance component estimates in random effect settings.
The ARC method offers a significant computational saving compared to ordinary maximum likelihood estimation for two reasons. In addition to the simplified integration, the reparameterization used by the ARC method conveniently separates estimation of the parameters: regression parameters are estimated from a standard probit model, whereas estimation of the dependence parameters require two distinct univariate optimizations of the row- and column-wise likelihoods.
In the next Theorem, we prove the weak consistency of the maximizer $\hat{\bm{\gamma}}$ of the ‘all’ likelihood (8) to the value $\bm{\gamma}$ of equation (7). First we introduce some notation. Let $Z_{ij}=1$ if $(\bm{x}_{ij},Y_{ij})$ is observed and take $Z_{ij}=0$ otherwise. The number of observations in row $i$ is $N_{i\text{\tiny$\bullet$}}=\sum_{j=1}^{C}Z_{ij}$ and similarly column $j$ has $N_{\text{\tiny$\bullet$}j}=\sum_{i=1}^{R}Z_{ij}$ observations in it. Let $\epsilon_{R}=\max_{i}N_{i\text{\tiny$\bullet$}}/N$ and $\epsilon_{C}=\max_{j}N_{\text{\tiny$\bullet$}j}/N$. In our motivating settings it is reasonable to assume that $\max(\epsilon_{R},\epsilon_{C})\to 0$ as $N\to\infty$. When we need a double sum over all pairs of data we use index $r$ as a ‘second $i$’ and $s$ as a ‘second $j$’. We take $Z_{ij}$ to be deterministic with at most one observation for any $(i,j)$ pair. In our motivating applications that would either never happen or be very rare. If rare, one could use the most recent observation only. The $\bm{x}_{ij}$ are not dependent on the $Z_{ij}$. Finally, $Y_{ij}$ are sampled from their probit distribution conditionally on $\bm{x}_{ij}$.
Theorem 1.
Let $Y_{ij}\in\{0,1\}$ follow the crossed random effects probit model (1) with true value $\bm{\gamma}_{0}$ for the parameter $\bm{\gamma}=\bm{\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ Let the number of observations $N\to\infty$ while $\max(\epsilon_{R},\epsilon_{C})\to 0$. Let $\bm{x}_{ij}\in\mathbb{R}^{p}$ satisfy
  1. 1.
    $\|\bm{x}_{ij}\|\leq B<\infty;$
  2. 2.
    $N^{-1}\sum_{ij}Z_{ij}\bm{x}_{ij}\bm{x}_{ij}^{\mathsf{T}}\to V\in\mathbb{R}^{p\times p},$ where $V$ is positive definite;
  3. 3.
    there is no nonzero vector $\bm{v}\in\mathbb{R}^{p}$ such that $\bm{v}^{\mathsf{T}}\bm{x}_{ij}\geq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=1$ and $\bm{v}^{\mathsf{T}}\bm{x}_{ij}\leq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=0$.
Let $\hat{\bm{\gamma}}\in\mathbb{R}^{p}$ be any maximizer of (8). Then for any $\epsilon>0$
$\Pr(\|\hat{\bm{\gamma}}-\bm{\gamma}_{0}\|>\epsilon)\to 0,\;\text{as $N\to\infty$}.$
The proof is reported in Section S1 of the Supplementary Material and we obtained it by adapting the proof strategy of Lumley and Mayer Hamblett (2003) to our setting.
Now we consider consistent estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the row and column likelihoods, respectively. Those provide consistent estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ with which one can then adjust the consistent estimate of $\bm{\gamma}$ to get a consistent estimate of $\bm{\beta}$.
Theorem 2.
Under the assumptions of Theorem 1, there is a root of the row likelihood equation that is a consistent estimator for $\tau^{2}_{A}$ and a root of the column likelihood equation that is a consistent estimator for $\tau^{2}_{B}.$
Our proof of this theorem is also reported in Section S1 of the Supplementary Material. The proof uses the subset argument of Jiang (2013) to show Cramer consistency of the MLE row likelihood and equivalently of the column likelihood. If one does not find Cramer consistency sufficient, then it is possible to construct a consistent estimator using all of the data from a subset of rows with a large number of observations. In Section S1 of the Supplementary Material we show how to construct such a consistent estimator and we give conditions under which the number of large rows will diverge to infinity as $N\to\infty$. See Theorem 3 of the Supplementary Material. We prefer our ARC estimator to an approach using just large rows, because it would be awkward to have to decide in practice which rows to use, and because we believe that there is valuable information in the other, smaller, rows.
3.2 Robust sandwich variance
After a customary Taylor approximation, the variance of $\hat{\bm{\theta}}$ is $\mathrm{var}(\hat{\bm{\theta}})\doteq D\,\mathrm{var}(\hat{\bm{\psi}})\,D^{\mathsf{T}},$ with $D$ the Jacobian matrix of the reparameterization from $\bm{\theta}$ to $\bm{\psi}$. Letting $\bm{u}_{\text{arc}}(\bm{\psi})$ denote the score vector of the ARC estimator constructed by stacking the scores of the three misspecified likelihoods, the asymptotic variance of $\hat{\bm{\psi}}$ is
$\mathrm{avar}(\hat{\bm{\psi}})=\mathcal{I}_{\text{arc}}^{-1}(\bm{\psi}_{0})V_{\text{arc}}(\bm{\psi}_{0})\mathcal{I}_{\text{arc}}^{-1}(\bm{\psi}_{0}),$ (12)
where $\mathcal{I}_{\text{arc}}({\bm{\psi}})=-\mathrm{E}(\partial\bm{u}_{\text{arc}}({\bm{\psi}})/\partial{\bm{\psi}})$ and $V_{\text{arc}}({\bm{\psi}})=\mathrm{var}(\bm{u}_{\text{arc}}({\bm{\psi}}))$ are the expected information and the score variance for the ARC estimator. These two matrices are not equal because the second Bartlett identity does not hold for misspecified likelihoods that constitute the ARC method. While estimation of the ‘bread’ matrix $\mathcal{I}_{\text{arc}}$ of the sandwich is not problematic, direct computation of the ‘filling‘ matrix $V_{\text{arc}}$ is not feasible in our large-scale setup because it requires us to approximate a large number of multidimensional integrals with a cost that does not meet our $\mathcal{O}(N)$ constraint.
Since ARC estimates require $\mathcal{O}(N)$ computations, we can estimate the variance of $\hat{\bm{\theta}}$ with a parametric bootstrap. However, it is preferable to evaluate the estimation uncertainty without assuming the correctness of the fitted model, for example using the nonparametric pigeonhole bootstrap described in Owen (2007). In that approach, the rows in the data set are resampled independently of the columns. So if a row is included twice and a column is included three times, the corresponding element is included six times. The resulting bootstrap variance for a mean (such as one in a score equation) over-estimates the random effects variance by an asymptotically negligible amount. It does not require homoscedasticity of either the random effects or the errors.
Now we describe a convenient approach that we have developed to estimate the variance of $\hat{\bm{\beta}}$ in $\mathcal{O}(N)$ operations without the need of resampling and repeated fitting as in the bootstraps mentioned above. The partitioned expected ARC information matrix is
$\mathcal{I}_{\text{arc}}(\bm{\psi})$ $=-\begin{pmatrix}\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial\bm{\gamma}\partial\bm{\gamma}^{\mathsf{T}})&0&0\\ \mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\bm{\gamma}\partial\tau^{2}_{A})&\mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\tau^{2}_{A}\partial\tau^{2}_{A})&0\\ \mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\bm{\gamma}\partial\tau^{2}_{B})&0&\mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\tau^{2}_{B}\partial\tau^{2}_{B})\\ \end{pmatrix}.$
Since the expected ARC information is triangular, then the asymptotic variance for $\hat{\bm{\gamma}}$ is
$\mathrm{avar}(\hat{\bm{\gamma}})=\mathcal{I}_{\text{all}}^{-1}(\bm{\gamma}_{0})V_{\text{all}}(\bm{\gamma}_{0})\mathcal{I}_{\text{all}}^{-1}(\bm{\gamma}_{0}),$ (13)
where $\mathcal{I}_{\text{all}}({\bm{\gamma}})=-\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial\bm{\gamma}\partial\bm{\gamma}^{\mathsf{T}})$ and $V_{\text{all}}({\bm{\gamma}})=\mathrm{var}(\partial\ell_{\text{all}}/\partial\bm{\gamma})$ are the Fisher expected information and the score variance of the marginal probit model. The asymptotic variance (13) for $\hat{\bm{\gamma}}$ is thus the same as that of the estimator that maximizes the all likelihood when the nuisance parameters $\tau^{2}_{A}$ and $\tau^{2}_{B}$ are known. The robust sandwich estimator of the variance of $\hat{\bm{\gamma}}$ is obtained by replacing $\mathcal{I}_{\text{all}}(\bm{\gamma}_{0})$ and $V_{\text{all}}(\bm{\gamma}_{0})$ with some estimators that are consistent and robust to misspecification. The expected information is naturally estimated with the observed information,
$J_{\text{all}}(\hat{\bm{\gamma}})=\sum_{ij}Z_{ij}\varphi(\hat{\eta}_{ij})\left(y_{ij}\frac{\varphi(\hat{\eta}_{ij})+\hat{\eta}_{ij}\,\Phi(\hat{\eta}_{ij})}{\Phi(\hat{\eta}_{ij})^{2}}+(1-y_{ij})\frac{\varphi(\hat{\eta}_{ij})-\hat{\eta}_{ij}\,\Phi(-\hat{\eta}_{ij})}{\Phi(-\hat{\eta}_{ij})^{2}}\right)\bm{x}_{ij}\bm{x}_{ij}^{\mathsf{T}},$
where $\hat{\eta}_{ij}=\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}.$ Estimation of $V_{\text{all}}(\bm{\gamma}_{0})$ is more involved. This matrix can be decomposed into the sum of three terms,
$V_{\text{all}}(\bm{\gamma}_{0})$ $=\sum_{ij}\sum_{rs}Z_{ij}Z_{rs}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{rs}^{\mathsf{T}}(\bm{\gamma}_{0})\right)$
$=\sum_{ijs}Z_{ij}Z_{is}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{is}^{\mathsf{T}}(\bm{\gamma}_{0})\right)+\sum_{ijr}Z_{ij}Z_{rj}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{rj}^{\mathsf{T}}(\bm{\gamma}_{0})\right)$
$-\sum_{ij}Z_{ij}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{ij}^{\mathsf{T}}(\bm{\gamma}_{0})\right),$
where $\bm{u}_{ij}(\bm{\gamma})$ is the score for a single observation $Y_{ij},$
$\bm{u}_{ij}(\bm{\gamma})=\frac{\varphi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\left(y_{ij}-\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\right)\bm{x}_{ij}}{\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})}.$
The corresponding estimator of $V_{\text{all}}(\bm{\gamma}_{0})$ is
$\widehat{V}_{\text{all}}(\hat{\bm{\gamma}})=\widehat{V}_{A}(\hat{\bm{\gamma}})+\widehat{V}_{B}(\hat{\bm{\gamma}})-\widehat{V}_{A\cap B}(\hat{\bm{\gamma}}),$
whose components are computed grouping the individual scores with respect to each random effect and their interaction,
$\widehat{V}_{A}(\hat{\bm{\gamma}})=\sum_{i}\bm{u}_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}})\bm{u}_{i\text{\tiny$\bullet$}}^{\mathsf{T}}(\hat{\bm{\gamma}}),\quad\widehat{V}_{B}(\hat{\bm{\gamma}})=\sum_{j}\bm{u}_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}})\bm{u}_{\text{\tiny$\bullet$}j}^{\mathsf{T}}(\hat{\bm{\gamma}}),\quad\widehat{V}_{A\cap B}(\hat{\bm{\gamma}})=\sum_{ij}Z_{ij}\bm{u}_{ij}(\hat{\bm{\gamma}})\bm{u}_{ij}^{\mathsf{T}}(\hat{\bm{\gamma}}),$
where $\bm{u}_{i\text{\tiny$\bullet$}}(\bm{\gamma})=\sum_{j}Z_{ij}\bm{u}_{ij}(\bm{\gamma})$ and $\bm{u}_{\text{\tiny$\bullet$}j}(\bm{\gamma})=\sum_{i}Z_{ij}\bm{u}_{ij}(\bm{\gamma}).$ Estimators of the form $\widehat{V}_{\text{all}}(\hat{\bm{\gamma}})$ are used in statistical modeling of data clustered within multiple levels in medical applications (Miglioretti and Heagerty, 2004) and in economics (Cameron et al., 2011) where they are known as two-way cluster-robust sandwich estimators.
Finally, we approximate the variance of $\hat{\bm{\beta}}$ by plugging in the estimates $\widehat{\sigma}_{A}^{2}$ and $\widehat{\sigma}_{B}^{2}$,
$\widehat{\mathrm{var}}(\hat{\bm{\beta}})=(1+\widehat{\sigma}_{A}^{2}+\widehat{\sigma}_{B}^{2})\,\widehat{\mathrm{var}}(\hat{\bm{\gamma}})=(1+\widehat{\sigma}_{A}^{2}+\widehat{\sigma}_{B}^{2})\,{J}_{\text{all}}^{-1}(\hat{\bm{\gamma}})\widehat{V}_{\text{all}}(\hat{\bm{\gamma}}){J}_{\text{all}}^{-1}(\hat{\bm{\gamma}}).$ (14)
A limitation of this approach is that it neglects the uncertainty in the estimation of the variance components: Although we do not expect a substantial impact in high-dimensional applications, it could be possible to adjust the (14) for the variability of the variance components through bootstrapping the row-wise and column-wise estimates $\hat{\tau}^{2}_{A}$ and $\hat{\tau}^{2}_{B}$.
4 Simulations
In this section, we simulate from the probit model with crossed random effects (1) and compare the performance of the ARC estimator with the traditional estimator obtained by maximizing the first-order Laplace approximation of the likelihood. A further method considered is given by an infeasible oracle estimator that uses the unknown true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to estimate the regression parameters as $\hat{\bm{\beta}}_{\mathrm{oracle}}=\hat{\bm{\gamma}}(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ The ARC method instead corrects $\hat{\bm{\gamma}}$ using estimates of the variance components. All methods were implemented in the R language. The package TMB (Kristensen et al., 2016) was used for the Laplace approximation, with the nlminb optimization function employed for its maximization. We did not use the popular glmer function from the R package lme4 (Bates et al., 2015) because TMB is substantially more computationally efficient and thus allowed us to compare our method to the first-order Laplace approximation at larger dimensions than it would otherwise be possible with glmer. The row- and column-wise likelihoods were coded in C++ and integrated in R with Rcpp (Eddelbuettel, 2013), and optimized by Brent’s method as implemented in the optimise function. The code is available in the public repository https://github.com/rugbel/arcProbit.
4.1 Simulation settings
We considered eight different settings defined by combining three binary factors. The first factor is whether the simulation is balanced (equal numbers of rows and columns) or imbalanced (with very unequal numbers) like we typically see in applications. The second factor is whether the regression model is null apart from a nonzero intercept or has nonzero regression coefficients. The third factor is whether the random effect variances are set at a high level or at a low level. Given $R$ and $C$, the set $\mathcal{S}$ is obtained by IID Bernoulli sampling with probability $N/(RC)$. This makes the attained value of $N$ random but with a very small coefficient of variation.
We denote by $R=N^{\rho}$ and $C=N^{\kappa}$ the number of rows and columns in the data, as a power of the total sample size $N$. The two levels of the balance factor are
 Balanced
$\rho=\kappa=0.56$, and
 Imbalanced
$\rho=0.88$ and $\kappa=0.53$.
This balanced case was already considered in Ghosh et al. (2022b). Because $\rho+\kappa>1$, the fraction of possible observations in the data is $N/(RC)=N^{1-\rho-\kappa}\to 0$ as $N\to\infty$ providing asymptotic sparsity in both cases. While the first choice has $R/C$ constant, the second choice has $R/C\to\infty$ with $N$. We believe that this asymptote is a better description of our motivating problems than either a setting with $C$ fixed as $R\to\infty$ or the setting common in random matrix theory (Edelman and Rao, 2005) where $R$ and $C$ diverge with $R/C$ approaching a constant value.
We considered seven predictors generated from a multivariate zero-mean normal distribution with a covariance matrix $\Sigma$ corresponding to an autocorrelation process of order one, so that the entry $(k,l)$ of $\Sigma$ is $\phi^{|k-l|}$. We set $\phi=0.5$ in all the simulations. We always used the intercept $\beta_{0}=-1.2$ because in our applications $\Pr(Y=1)<1/2$ is typical. For the predictor coefficients we considered the following two choices
 Null
$\beta_{\ell}=0$ for $\ell=1,\ldots 7$, and
 Linear
$\beta_{\ell}=-1.2+0.3\ell$ for $\ell=1,\dots,7.$
The first setting is a null one where $\bm{x}$ is not predictive at all, while the second setting has modestly important nonzero predictors whose values are in linear progression.
The two choices for the variance component parameters are
 High variance
$\sigma_{A}=1$ and $\sigma_{B}=1$, and
 Low variance
$\sigma_{A}=0.5$ and $\sigma_{B}=0.2$.
We chose the first setting to include variances higher than what is typically observed in applications. The second setting is closer to what we have seen in data such as that in Section 5. We represent the eight settings with mnemonics as shown in Table 1. For example, Imb-Nul-Hi  means row-column imbalance ($\rho=0.88,\kappa=0.53$), all predictor coefficients are zero and the main effect variances are large ($\sigma_{A}=1,\sigma_{B}=1$).
Setting Sparsity Predictors Variances
$\rho$ $\kappa$ $\sigma_{A}$ $\sigma_{B}$
Bal-Nul-Hi 0.56 0.56 all zero 1.0 1.0
Imb-Nul-Hi 0.88 0.53 all zero 1.0 1.0
Bal-Lin-Hi 0.56 0.56 not all zero 1.0 1.0
Imb-Lin-Hi 0.88 0.53 not all zero 1.0 1.0
Bal-Nul-Lo 0.56 0.56 all zero 0.5 0.2
Imb-Nul-Lo 0.88 0.53 all zero 0.5 0.2
Bal-Lin-Lo 0.56 0.56 not all zero 0.5 0.2
Imb-Lin-Lo 0.88 0.53 not all zero 0.5 0.2
Table 1: Summary of the eight simulation settings.
For each of these eight settings, we considered 13 increasing sample sizes $N$ in the interval from $10^{3}$ to $10^{6}$ obtained by taking 13 equispaced values on the $\log_{10}$ scale. As we see next, the Laplace method had a cost that grew superlinearly and to keep costs reasonable we only used sample sizes up to $10^{5}$ for that method. For each of these 13 sample sizes and each of the eight settings, we simulated 1000 data sets.
Graphs comparing the computational costs, the statistical properties and the scalability of the three estimation methods for all eight settings are reported in the Supplementary Materials. To save space, we present graphs for only one of the settings, Imb-Nul-Hi, and just summarize the other settings. This chosen setting is a challenging one. It is not surprising that imbalance and large variances are challenging. The binary regression setting is different from linear modeling where estimation difficulty is unrelated to predictor coefficient values. The main reason to highlight this setting is that it illustrates an especially bad outcome for the Laplace method estimate of $\sigma_{A}$. Similar but less extreme difficulties for the Laplace method’s estimates of $\sigma_{A}$ appear in setting Imb-Nul-Lo.
4.2 Computational cost
Computational cost is the first and most critical quantity to consider since methods that are too expensive cannot be used, regardless of their quality. Figure 1 shows the average computation times for the three methods, obtained on a 16-core 3.5 GHz AMD processor equipped with 128 GB of RAM. It also shows regression lines of log cost versus $\log(N)$, marked with the regression slopes. The ARC and oracle method’s slope are both very nearly $1$ as expected. The Laplace method slope is clearly larger than $1$ and, as noted above, we curtailed the sample sizes used for that case to keep costs reasonable. Similar estimated computational costs were obtained for the other seven settings as shown in the Supplementary Material are summarized in Table 2 which shows estimated computational cost rates for all eight settings. The ARC slope is consistently very close to $1$. The oracle slope is consistently close to $1$ in imbalanced settings and somewhat bigger than $1$ in balanced ones. The Laplace slope is consistently larger than $1$. It tends to be higher for balanced settings although one of the imbalanced cases also has a large slope.
We have investigated the data behind the oracle slopes, but cannot yet explain the mild superlinearity we sometimes see. The number of Fisher-scoring iterations used by the oracle method varies at small sample sizes but is consistently near seven at larger sample sizes where the superlinearity is more prominent. We have seen some outliers in the computation times at small sample sizes but not at large sample sizes. Replacing the means at different $N$ by medians does not materially change the slopes. We consider the amount of unexplained nonlinearity to be small but not negligible. For example, when going from $N=10^{3}$ to $N=10^{6}$ a rate like $N^{1.06}$ yields a roughly 1500 fold cost increase instead of the expected 1000 fold increase. The anomaly is concentrated in the balanced simulations, but knowing this has not been enough to identify the cause.
Setting Oracle slope ARC slope Laplace slope
Bal-Nul-Hi 1.06 1.01 1.30
Bal-Nul-Lo 1.05 1.01 1.24
Bal-Lin-Hi 1.06 0.99 1.29
Bal-Lin-Lo 1.07 0.99 1.26
Imb-Nul-Hi 0.98 0.99 1.23
Imb-Nul-Lo 0.98 1.00 1.13
Imb-Lin-Hi 1.02 0.99 1.15
Imb-Lin-Lo 0.98 0.99 1.15
Table 2: For all eight settings, this table shows the observed computational cost as a rate in $N$, fit by linear regression of the logarithm of average CPU seconds on $\log(N)$.
Figure 1: Time in seconds versus sample size N for the Imb-Nul-Hi setting for all three algorithms: Laplace, ARC and the Oracle.
4.3 Regression coefficient estimation
Next, we turn to estimation of the regression coefficients, treating the intercept differently from the others. The intercept poses a challenge because it is somewhat confounded with the random effects. For instance, if we replace $\beta_{0}$ by $\beta_{0}+\lambda$ while replacing $a_{i}$ by $a_{i}-\lambda$ then $Y_{ij}$ are unchanged. Large $\lambda$ would change $\bar{a}=(1/R)\sum_{i=1}^{R}a_{i}$ by an implausible amount that should be statistically detectable given that $a_{i}\sim\mathcal{N}(0,\sigma^{2}_{A})$. On the other hand $|\lambda|=O(\sigma_{A}/\sqrt{R})$ would be hard to detect statistically. The other regression parameters are not similarly confounded with main effects in our settings. We note that when a predictor is a categorical variable that is either a function of the row index $i$ or of the column index $j$, then it can be subject to some confounding similar to the intercept’s confounding. This was observed in Ghosh et al. (2022b).
Because of the confounding described above, we anticipate that the true MSE rate for the intercept cannot be better than $\mathcal{O}(\min\{R,C\}))$ which is $\mathcal{O}(N^{-0.53})$ in our imbalanced settings and $\mathcal{O}(N^{-0.56})$ in our balanced settings. For the other coefficients $\mathcal{O}(N^{-1})$ is not ruled out by this argument. Figures 2 and 3 reports the mean square errors for the intercept and the coefficient of the first predictor estimated for different sample sizes, for the three estimators under study in the Imb-Nul-Hi setting. We report here the plot only for the first predictor because the mean squared errors of the estimates of the seven regression coefficients were essentially equivalent. All three estimators show an MSE very close to $\mathcal{O}(N^{-1})$ for $\beta_{1}$. Where we anticipated an MSE no better than $\mathcal{O}(N^{-0.53})$ (imbalanced) or $\mathcal{O}(N^{-0.56})$ (balanced) for the intercept we saw slightly better MSEs with slopes between $-0.57$ and $-0.60$, confirming our expectation that the intercept would be harder to estimate than the regression coefficients.
Figure 2: Mean squared errors for coefficient estimates in the Imb-Nul-Hi setting. The top row is for $\beta_{0}$ and the bottom row is for $\beta_{1}$. Each method is shown with a line and solid points and fitted reference lines for the other two methods (that overlap somewhat).
4.4 Variance component estimation
Here we present the estimation errors in the variance component parameters $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. The oracle method is given the true values of these parameters and so the comparison is only between ARC and the Laplace method. For the variance parameter $\sigma^{2}_{A}$, the data only have $R$ levels $a_{1},\dots,a_{R}$. If they were observed directly, then we could estimate $\sigma^{2}_{A}$ by $(1/R)\sum_{i=1}^{R}a_{i}^{2}$ and have an MSE of $\mathcal{O}(R^{-1})$. In practice, the $a_{i}$ are obscured by the presence of the signal, the noise $\varepsilon_{ij}$ and the other random effects $b_{j}$. Accordingly, the best rate we could expect for $\sigma^{2}_{A}$ is $\mathcal{O}(R^{-1})=\mathcal{O}(N^{-\rho})$ and the best we could expect for $\sigma^{2}_{B}$ is $\mathcal{O}(C^{-1})=\mathcal{O}(N^{-\kappa})$.
Figure 3 shows the MSE for estimation of $\sigma_{A}$ and $\sigma_{B}$ for ARC and Laplace methods in the Imb-Nul-Hi setting. Due to the imbalance, our anticipated rates are $\mathcal{O}(N^{-0.88})$ for $\sigma_{A}$ and $\mathcal{O}(N^{-0.53})$ for $\sigma_{B}$. The ARC method does slightly better than these rates. The Laplace method attains nearly this predicted rate for $\sigma_{B}$ but does much worse for $\sigma_{A}$. We can understand both of these discrepancies in terms of biases, described next. Figure 4 shows boxplots for the parameter estimates of $\sigma_{A}$ and $\sigma_{B}$ with the ARC and Laplace methods. We can compare the center of those boxplots to the reference line at the true parameter values and see that for the ARC estimates there is a bias decreasing at a faster rate than the width of the boxes. This explains the slightly better than predicted rates that we see for ARC. Instead the Laplace method has a substantial bias that only decreases very slowly as $N$ increases, giving Laplace a worse than expected rate, especially for $\sigma_{A}$.
In order to understand the difficulties of the Laplace approximation method we turn to the literature on Laplace approximations where the dimension of the integral grows with the sample size. The crossed random effects integral has dimension $R+C$ for $N$ observations. The largest number of parameters in any of our simulations was for the imbalanced setting with $R+C=\mathcal{O}(N^{0.88})$ and this is where the Laplace approximation had its worst performance. Section 6 of Shun and McCullagh (1995) suggests that a sufficient condition for the Laplace approximation to be reliable is, in our notation, that $R+C=o(N^{1/3})$. This is not possible in a sparse setting with $RC\gg N$. Ogden (2021) notes that Shun and McCullagh (1995) use an assumption on the growth rate of the derivatives of the exponent in the Laplacian integrand and says that assumption is not realistic for generalized linear mixed models. She then gives an analysis of the hierarchical setting which we believe is much simpler than the present crossed setting. For a hierarchical model with our row random effects but no column random effects, the likelihood is a product of $R$ one dimensional integrals. The logarithm of the $i$’th integrand is a sum of $N_{i\text{\tiny$\bullet$}}$ terms. She notes that the error in a first order Laplace approximation is $\mathcal{O}(\sum_{i}N_{i\text{\tiny$\bullet$}}^{-1})$. The smallest that could be with $N$ observations in $R$ columns is $R(N/R)^{-1}=R^{2}/N$, a quantity that is $\mathcal{O}(N^{2\rho-1})=\mathcal{O}(N^{0.76})$ in our unbalanced setting. Things improve with higher order approximations. For the definition of $k$’th order in Ogden (2021), equation (10) there gives an error bound of $\mathcal{O}(\sum_{i}N_{i\text{\tiny$\bullet$}}^{-k})$. Ogden (2021) remarks that this bound might be conservative. All of our simulations had $\sum_{i}N_{i\text{\tiny$\bullet$}}^{-1}$ increasing with $N$ but only in the imbalanced null cases did it look like the Laplace method might be inconsistent (due to bias). For the approximation of one-dimensional integrals in ARC an accurate numerical approximation was employed and we saw empirical convergence rates close to those we anticipated.
Figure 3: MSEs and estimated convergence rates for $\sigma_{A}$ and $\sigma_{B}$ for ARC and Laplace methods in setting Imb-Nul-Hi. In each plot there is a reference line for the covergence rate of the other method.
Figure 4: Boxplots of $\sigma_{A}$ and $\sigma_{B}$ estimates for the ARC and Laplace methods in the setting Imb-Nul-Hi. There are horizontal reference lines at the true parameter values. The ARC estimates have negative bias at small $N$ while the Laplace estimates have a strong extreme negative bias for all $N$.
4.5 Other settings
The simulation results for all eight settings, are reported in full in the Supplementary Figures S5–S44. We already discussed computational cost in eight settings based on Table 2. Here we make brief accuracy comparisons. We compare our proposed ARC method to the oracle method which is infeasible because it requires knowledge of $\sigma_{A}$ and $\sigma_{B}$ and to the Laplace method which becomes infeasible for large $N$ because it does not scale as $\mathcal{O}(N)$. In most settings and for most parameters, the oracle method was slightly more accurate than ARC. That did not always hold. For instance in setting Bal-Lin-Hi (Figure S14) we see that ARC was slightly more accurate than the oracle method for $\beta_{1}$. The figures in the Supplementary material show some cases where ARC has a slight advantage over the Laplace method and some where it has a slight disadvantage. There are a small number of cases where the Laplace method has outliers at $N=10^{3}$ that cause the linear regression slope to be questionable. These are present in setting Imb-Lin-Hi (Figure S20). However the attained MSEs at $N=10^{5}$ do not differ much between ARC and Laplace in that setting. Our conclusion is that compared to the Laplace method, ARC is scalable and robust.
The Supplementary Material Section S2 makes a comparison of ARC with the maximum pairwise likelihood estimator of Bellio and Varin (2005). The comparison is made for the Imb-Nul-Hi setting that we have been focusing on. From Section 2, the cost of pairwise composite likelihood is $\Omega(\max(N^{2}/R,N^{2}/C))=\Omega(\max(N^{2-\rho},N^{2-\kappa}))$. For the imbalanced setting this is $\Omega(N^{2-0.53})=\Omega(N^{1.47}).$ We see in Section S2 that the empirical cost of the pairwise composite likelihood grows as $N^{1.42}$, close to the predicted rate. The pairwise method attains very similar parameter estimation accuracy to ARC. The most important difference for this example is that ARC costs $\mathcal{O}(N)$ while pairwise likelihood is far more expensive and does not scale to large data sets.
5 Application to the Stitch Fix Data
In this section we illustrate our ARC method on a data set from Stitch Fix. Quoting from Ghosh et al. (2022b):
Stitch Fix is an online personal styling service. One of their business models involves sending customers a sample of clothing items. The customer may keep and purchase any of those items and return the others. They have provided us with some of their client ratings data. That data was anonymized, void of personally identifying information, and as a sample it does not reflect their total numbers of clients or items at the time they provided it. It is also from 2015. While it does not describe their current business, it is a valuable data set for illustrative purposes.
The Stitch Fix data consists of $N=5{,}000{,}000$ ratings from $R=744{,}482$ clients on $C=3{,}547$ items. The data also includes client- and item-specific covariates. In this data, the binary response $Y_{ij}$ of interest was whether customer $i$ thought that item $j$ was a top rated fit or not, with $Y_{ij}=1$ for an answer of ‘yes’. The predictor variables we used are listed in Table 3. Some of the categorical variable levels in the data had only a small number of levels. The table shows how we have aggregated them.
Variable Description Levels
Client fit client fit profile fitted
loose or oversize
straight or tight
Edgy edgy style? yes/no
Boho Bohemian style? yes/no
Chest chest size numeric
Size dress size numeric
Material item primary material artificial fiber
leather or animal fiber
regenerated fiber
vegetable fiber
Item fit item fit fitted
loose or oversized
straight or tight
Table 3: Predictors available in the Stitch Fix data. There is one block of client predictors followed by a block of item predictors.
As anticipated in Section 2, we assume missing data at random and have removed rows with any missing entries. That leaves us with $N=4{,}965{,}960$ of the original $5{,}000{,}000$ ratings. We have data from $741{,}221$ clients on $3{,}523$ items. The data are not dominated by a single row or column. The customer with the most records accounts for $N\epsilon_{R}$ records where $\epsilon_{R}\doteq 1.25\times 10^{-5}$. The item with the most records accounts for $N\epsilon_{C}$ of them with $\epsilon_{C}\doteq 2.77\times 10^{-2}$. The data are sparse because $N/(RC)\doteq 1.9\times 10^{-3}.$
In a business setting one would fit and compare a wide variety of different binary regression models in order to understand the data. Our purpose here is to study large scale probit models with including crossed random effects and so we choose just one model for illustration, possibly the first model of many that one would consider. We consider the probit model with crossed random effects whose fixed effects are specified according to the symbolic model formula:
Top $\sim$ Client fit + Edgy + Boho + Chest + Size + Material + Item fit,
where Top is the binary response variable described earlier. The model has $p=12$ parameters for fixed effects including the intercept. The first level of each categorical predictor in alphabetical order (Table 3) is used as the reference level in fitting the model.
Table 4 reports 1) the maximum likelihood estimates of the regression parameters under a marginal probit model that ignores the customer and item heterogeneity, and 2) the ARC estimates for the probit model with two crossed random effects for the customers and items. The z values reported in the table are computed with the observed information for the marginal probit model and with the two-way cluster-robust sandwich estimator described in Section 3.2. The ARC estimates of the variance components are $\hat{\sigma}_{A}=0.53$ and $\hat{\sigma}_{B}=0.34$. As expected, ignoring the customer and item heterogeneity lead to a large underestimation of the uncertainty in the parameter estimates and thus in the marginal probit all the predictors are strongly significant given the very large sample size. Conversely, the cross-random effects model takes into account the sources of heterogeneity and leads us to learn that the item fit is not a significant predictor of the top rank and that items made by vegetable fibers are less likely to be ranked top compared to clothes made by artificial fibers.
While we have emphasized the naive variance estimation of the marginal probit model, Table 4 also shows the attenuation bias. The naive model systematically underestimates the absolute value of the coefficients while simultaneously overestimating their statistical significance.
Marginal probit Random-effects probit
Variable ${}^{\text{\textdagger}}$ z value p value ${}^{\text{\textdagger}}$ z value p value
[…]
Edgy yes --3.0 --25.08 $<0.001$-3.5 --7.49 $<0.001$
[…]
Chest --0.5 --12.30 $<0.001$-0.6 --7.32 $<0.001$
[…]
Material leather or animal --12.9 --12.99 $<0.001$-15.2 --1.57 0.116
[…]
vegetable --12.2 --58.39 $<0.001$-14.5 --3.13 $0.002$
Item fit loose or oversized  9.7 36.15 $<0.001$   11.4 1.78 0.075
straight or tight $-2.1$-9.55 $<0.001$-2.5 --0.67 0.500
Table 4: Stitch Fix Binary Regression Results. ${}^{\text{\textdagger}}$ll the estimated predictor parameters are multiplied by 100. Nonsignificant p values are marked in bold. See Table 3 for a description of the predictors. The first level of each categorial predictor in alphabetic order is used as reference level.
Figure 5 compares the two-way cluster robust sandwich standard errors with 1) the naive standard errors from a probit fit ignoring dependence between items and customers and 2) the pigeonhole nonparametric bootstrap standard errors of Owen (2007) that do not assume correct model specification, as mentioned in Section 3.2. The figure shows how closely the sandwich and pigeonhole standard errors agree, while the naive standard errors are much smaller. The naive values correspond to variances underestimated by factors ranging from 4 to just over 1240, depending on the parameter and only slightly on whether we use sandwich or pigeonhole estimates of the coefficient variances. These standard errors are reported in Table S1 of the Supplementary Material. We also computed parametric bootstrap standard errors (not shown). Those were in between the naive ones and the nonparametric ones.
Figure 5: Stitch Fix Application. Comparison between the two-way cluster robust sandwich standard errors (x axis), the naive GLM standard errors (y axis, triangles) and the Pigeonhole nonparametric bootstrap standard errors (y axis, circles). The gray reference line is the first quadrant bisector $y=x$. All standard errors are multiplied by 100.
6 Bibliographic Notes
In this section we review some related literature on the difficulties that crossed random effects bring. McCullagh (2000) points out a key difference between the crossed and nested random effects. Even in an intercept-only setting with $Y_{ij}=\mu+a_{i}+b_{j}+\varepsilon_{ij}$ and all $R\times C$ values observed, there is no correct bootstrap for the variance of the simple average of the $Y_{ij}$. Resampling rows and columns independently is mildly conservative in this case. That mild conservatism also extends to some sparse settings (Owen, 2007). Gao and Owen (2017) provide an $\mathcal{O}(N)$ algorithm to consistently estimate an intercept plus variance components. Gao and Owen (2020) provides an $\mathcal{O}(N)$ algorithm for linear regression with crossed random effects; that algorithm is less efficient than generalized least squares but the variance estimates from it properly account for all of the random effects. They also show that computing the Gaussian likelihood for crossed random effects at just one parameter value costs $\Omega(N^{3/2})$.
Ghosh et al. (2022a) provide a backfitting algorithm for generalized least squares that costs $\mathcal{O}(N)$ per iteration and they give asymptotic conditions under which the number of iterations is $\mathcal{O}(1)$. Ghandwani et al. (2023) extended the algorithm to a linear mixed model that include random slopes. Ghosh et al. (2022b) consider logistic regression with random intercepts adapting an integral approximation from Schall (1991). Their iterations cost $\mathcal{O}(N)$ and empirically the algorithm runs in $\mathcal{O}(N)$ time, but there is no proven bound on the number of iterations required.
The computational challenge from crossed random effects is also present for Bayesian methods. A plain Gibbs sampler costs $\mathcal{O}(N)$ per iteration but takes $\Omega(N^{1/2})$ iterations for the model in Gao and Owen (2017). Papaspiliopoulos et al. (2023) derived a scalable MCMC algorithm for generalized linear models with crossed-random effects that extends previous work on a collapsed Gibbs sampler described in Papaspiliopoulos et al. (2020). This Bayesian simulation method is more general than our proposal because it is not restricted to a particular type of response. Our proposal instead derives a frequentist solution to the important problem of consistent and scalable estimation of a regression model for cross-classified binary responses. Differently from the MCMC method developed by Papaspiliopoulos et al. (2023), our approach does not require the use of simulation since it is only necessary to compute a large number of one-dimensional integrals which we approximate through an adaptive quadrature carefully implemented to avoid the risk of accumulating integration errors.
Other recent work has considered variational approximations to build computational efficient fitting of generalized linear mixed models with crossed-random effects. Goplerud et al. (2023) developed a variational approximation for scalable Bayesian estimation using an appropriate relaxing of the mean-field assumption to avoid underestimation of posterior uncertainty in high-dimensions. Parallel work on frequentist inference for generalized linear models with crossed-random effects has been conducted by Xu et al. (2023) by combining variational approximations and composite likelihoods that consider row-column decomposition in a similar way to ours. This approach is particularly convenient for Poisson and gamma regression models because in this case analytical calculations allow the approximation of one-dimensional integrals that appear in the composite likelihood to be avoided. However, the consistency of the Xu et al. (2023) estimation method for binary data has not been established yet, and only a conjecture for the logistic case has been formulated by the authors. Further investigation of the inferential properties of their proposal for the case of binary data seems warranted for a thorough comparison of the two methods. Hall et al. (2020) consider message passing algorithms for generalized linear mixed models and their Section 6 includes a crossed effecs binary regression. However it has $R=10$ and $C=6$ in our notation with $3$ replicates at each $(i,j)$ pair, and it did not address scalability.
Bartolucci et al. (2017) consider another composite likelihood that combines a row likelihood with a column likelihood to estimate a hidden Markov model for two-way data arrays. This approach shares the philosophy of our proposal but differs from our work in terms of model (latent discrete Markov variables vs crossed random effects), fitting procedure (EM algorithm vs direct maximization), data structure (balanced vs unbalanced and sparse) and motivating application (genomics vs e-commerce).
7 Conclusions
We have developed a method for crossed random effects for binary regression. It is a new version of composite likelihood that, by exploiting the connection between the Gaussian latent variables and the Gaussian random effects in the probit model, makes for very convenient factors in the composite likelihood.
In simulations, we find that the computational cost of our ARC proposal is $\mathcal{O}(N)$ and its accuracy is competitive with and even more robust than a computationally infeasible Laplace approximation that has a cost which grow superlinearly with $N$. The mean squared errors of our method are very nearly $\mathcal{O}(\min\{R,C\}^{-1})$ for the intercept and $\mathcal{O}(N^{-1})$ for the other regression parameters. They are nearly $\mathcal{O}(R^{-1})$ and $\mathcal{O}(C^{-1})$ for the row- and column-random effect variance parameters, respectively.
In the application to the e-commerce Stitch Fix data, a naive probit model finds that the coefficients of a model with 12 regression coefficients are all very statistically significant. After properly accounting for correlations due to random effects, we find that only eight of them remain significant at the 5% level. The naive analysis might lead to one following false leads. It can even lead one to confidently attribute incorrect signs to some parameters, even though $\mathrm{sign}(\gamma_{j})=\mathrm{sign}(\beta_{j})$. This can happen when $\mathrm{sign}(\hat{\gamma}_{j})\neq\mathrm{sign}(\gamma_{j})$ and the estimate of $\mathrm{var}(\hat{\gamma}_{j})$ is unrealistically small.
One surprise in this work was the difficulty of getting good approximations to even the one dimensional Gaussian integrals that arise. We are grateful for the insights in Ogden (2021) and Shun and McCullagh (1995). We surmise that the difficulties in multidimensional integration for random effects models on big data sets must be even more severe than the univariate ones we saw.
We have used a probit model instead of a logistic one because a Gaussian latent variable is a very natural counterpart to the Gaussian random effects that are the default in random effects models. That connection simplified our modeling and computation. As noted in the introduction, logit and probit link functions are nearly proportional outside of tail regions.
Gibbons and Hedeker (1997) remark that “As in the case of fixed effect models, selection of probit versus logistic response functions appears to have more to do with custom or practice within a particular discipline than differences in statistical properties.” Thus, while a logistic regression approach is outside the scope of the present paper we think that an extension to the logistic case is worth pursuing and that our present findings provide a good starting point. If we had used a logistic instead of the probit model, then the all-likelihood would have required us to approximate further $N$ univariate integrals with respect to a convolution of the Gaussian and logistic probability density functions. That would be much more expensive than the probit model costs, although still within the $\mathcal{O}(N)$ constraint.
Another approach to extending the ARC method to logistic regression uses the fact that the logistic distribution is a scale mixture of the standard normal distribution (Stefanski, 1991). This property is often used in Bayesian computations, see, for example, Narisetty et al. (2018).
Supplementary Material
The supplementary material contains proofs of the theorems of Section 3, additional simulation results discussed in Section 4, further references to the literature and a summary table of the binary regression standard errors for the Stitch Fix data.
R code for replicating our results is available in the public repository https://github.com/rugbel/arcProbit.
Acknowledgements
We thank Bradley Klingenberg and Stitch Fix for making this data set available. This work was supported in part by grants IIS-1837931 and DMS-2152780 from the U.S. National Science Foundation, IRIDE from Ca’ Foscari University of Venice, and by the research project Latent Variable Models for Complex Data founded by the European Union - NextGenerationEU (MUR DM funds 737/2021).

### v4

@@
-Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects
+Consistent and Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects
 Abstract
-Crossed random effects structures arise in many scientific contexts.
-They raise severe computational problems with likelihood computations scaling like $N^{3/2}$ or worse for $N$ data points.
-In this paper we develop a new composite likelihood approach for crossed random effects probit models.
-For data arranged in $R$ rows and $C$ columns, the likelihood function includes a very difficult $R+C$ dimensional integral.
-The composite likelihood we develop uses the marginal distribution of the response along with two hierarchical models.
-The cost is reduced to $\mathcal{O}(N)$ and it can be computed with $R+C$ one dimensional integrals.
-We find that the commonly used Laplace approximation has a cost that grows superlinearly.
-We get consistent estimates of the probit slope and variance components from our composite likelihood algorithm.
-We also show how to estimate the covariance of the estimated regression coefficients.
-The algorithm scales readily to a data set of five million observations from Stitch Fix with $R+C>700{,}000$.
-Keywords: Composite Likelihood; High-Dimensional Binary Data; Probit Regression; Scalability.
+Estimation of crossed random effects models commonly requires computational costs that grow faster than linearly in the sample size $N$, often as fast as $\Omega(N^{3/2})$, making them unsuitable for large data sets.
+For non-Gaussian responses, integrating out the random effects to get a marginal likelihood brings significant challenges, especially for high dimensional integrals where the Laplace approximation might not be accurate.
+We develop a composite likelihood approach to probit models that replaces the crossed random effects model with some hierarchical models that require only one-dimensional integrals.
+We show how to consistently estimate the crossed effects model parameters from the hierarchical model fits.We find that the computation scales linearly in the sample size.
+We illustrate the method on about five million observations from Stitch Fix where the crossed effects formulation would require an integral of dimension larger than $700{,}000$.
+Keywords: Adaptive Gauss-Hermite Quadrature; Binary Regression; E-Commerce; High-Dimensional Data
 1 Introduction
-In this paper we develop a scalable approach to mixed effects modeling with a probit link and a crossed random effects error structure.
-The common notation for mixed effects models incorporates the fixed and random effects through a formula such as $X\bm{\beta}+Z\bm{b}$ using matrices $X$ and $Z$ of known predictors and unknown coefficient $\bm{\beta}$ and $\bm{b}$ where $\bm{b}$ happens to be random.
-This formulation is simple and elegant but it hides an important practical difference.
-When the random effects have a hierarchical structure then computations usually scale well to large problems.
-However, when the random effects have a crossed structure, then computations can easily grow superlinearly, which greatly restricts the use of these models.
-For instance, computing the maximum likelihood estimator in a Gaussian model with crossed random effects by the usual algorithms includes linear algebra with a cost that grows at least as fast as $N^{3/2}$ given $N$ data points (Gao and Owen, 2020), while remaining $\mathcal{O}(N)$ in the hierarchical setting.
-The binary regression setting we consider here has the further difficulty of a very high dimensional integral that is not present for a model with a Gaussian response.
-Maximizing the likelihood with the expectation-maximization algorithm raises the same intractability problems as direct likelihood maximization because the integral appearing in the expectation step has the same very high dimension as the integral appearing in the likelihood.
-In an era where the size $N$ of data sets is growing rapidly it is not possible to use algorithms with a cost of $\Omega(N^{3/2})$.
-Ideally the cost should be $\mathcal{O}(N)$.
+In this paper we develop a new composite likelihood approach to handling probit models with two crossed random effects.
+Our initial motivation was to get point and interval parameter estimates at a computational cost that grows only linearly in the sample size, $N$.
+Standard algorithms for crossed random effects typically have superlinear cost, commonly $\Omega(N^{3/2})$, making them unsuitable for modern large data sets.
+A second issue is that the marginal likelihood in a crossed random effects model is an integral over $\mathbb{R}^{D}$ where $D$ is large enough to make the integration problem challenging.
+Our scalable estimation method replaces this $D$-dimensional integral by $D$ integrals of dimension one.
+The common notation for mixed effects models combines fixed and random effects through a formula such as $X{\beta}+Z{b}$ using matrices $X$ and $Z$ of known predictors with unknown coefficient vectors ${\beta}$ and ${b}$ where ${b}$ happens to be random.
+This formulation is simple and elegant but it hides some extremely important practical differences.
+As discussed above, the crossed setting leads to one high-dimensional integral while the hierarchical one uses many low-dimensional integrals.
+For Gaussian responses we can use generalized least squares on the response vector, without explicitly solving an integral.
+Even there, the crossed setting is harder.
+A hierarchical model has a block diagonal covariance matrix for the response vector resulting in a linear cost.
+For unbalanced crossed random effects, generalized least squares typically has a superlinear cost.
+With the size $N$ of data sets growing rapidly it is not possible to use estimation methods with a cost of $\Omega(N^{3/2})$.
+Ideally the cost should be $O(N)$.
 We are motivated by electronic commerce problems with large data sets.
 A company might have customers $i=1,\dots,R$, to which it sells items $j=1,\dots,C$.
-It would then be interested in modeling how a response $Y_{ij}$ depends on some predictors $\bm{x}_{ij}\in\mathbb{R}^{p}$.
-If they do not account for the fact that $Y_{ij}$ and $Y_{is}$ are correlated due to a common customer $i$ or that $Y_{ij}$ and $Y_{rj}$ are correlated due to a common item $j$, they will not get efficient estimates of the coefficients in their model.
-Much more seriously, they will not get reliable standard errors for those estimates.
-In a setting where accounting for random effects is computationally impossible we expect that many users will simply ignore them, getting very naive variance estimates and finding too many things significant.
-A typical feature of data in our motivating applications is a very sparse sampling.
-Only $N\ll RC$ of the possible $(\bm{x}_{ij},Y_{ij})$ values are observed.
+It would then be interested in modeling how a response $Y_{ij}$ depends on some predictors ${x}_{ij}\in\mathbb{R}^{p}$.
+If they do not account for the fact that $Y_{ij}$ and $Y_{is}$ are correlated due to a common customer $i$ or that $Y_{ij}$ and $Y_{rj}$ are correlated due to a common item $j$, they will get an inefficient estimate.
+This flaw is less severe when $N$ is so large.
+What is very serious is that they will get unreliable standard errors for their estimates.
+In a setting where accounting for random effects is computationally impossible we expect that many users will simply ignore them, getting very naïve variance estimates and finding too many things significant.
+A typical feature of data in our motivating applications is very sparse and imbalanced sampling.
+Only $N\ll RC$ of the possible $({x}_{ij},Y_{ij})$ values are observed.
 There is generally no simple structure in the pattern of which $(i,j)$ pairs are observed.
-While we use electronic commerce as our example, large sparsely sampled crossed data sets arise in other areas too.
-One of the main open questions in generalized linear mixed models, like the ones we consider here, is at what rate do the estimators of the model parameters converge?
-Intuitively one might expect $\mathcal{O}(N^{-1/2})$ or $\mathcal{O}(\min\{R,C\}^{-1/2})$.
-In some settings estimators for different parameters converge at different rates.
-See Jiang (2013) for some discussion.
-It was only comparatively recently that consistency of maximum likelihood estimation was proved in some generalized linear mixed models.
-This is also due to Jiang (2013).
-In this paper we focus on binary responses, $Y_{ij}\in\{0,1\}$.
-In commerce applications, such responses may be more important than real-valued ones.
-Examples include whether customer $i$ purchased item $j$, or returned it or reported satisfaction with it.
-In addition to the numerical algebra that underlies the $\Omega(N^{3/2})$ cost of standard algorithms this setting brings a second difficulty of an $R+C$ dimensional integral over random effect values in the likelihood.
-It is very common to consider a logistic regression for such binary responses.
-In this paper we consider a probit regression.
-The probit model is especially suitable because it is typical to use Gaussian models for the random effects.
+It is common for the data to have very unequal sampling frequencies in each of the row and column variables.
+The scaling problem is easiest to describe for generalised least squares solutions to linear mixed models, based on results from Gao & Owen (2020).
+They note that the algorithm for generalised least squares involves solving a system of $R+C$ equations in $R+C$ unknowns which has cost $\Omega\{(R+C)^{3}\}$ in standard implementations.
+If $RC>N$ then $\max(R,C)>N^{1/2}$ and so $(R+C)^{3}>N^{3/2}$.
+The average number of observations per level is $N/(RC)$.
+This ratio is well below one in our motivating applications, and as long as it is $o(N^{1/3})$, the cost of standard algebra will be superlinear.
+Standard Bayesian solutions run into a similar difficulty.
+For an intercept plus crossed random effects model, Gao & Owen (2017) show that the Gibbs sampler takes $\Omega(N^{1/2})$ iterations that each have a cost proportional to $N$, for $\Omega(N^{3/2})$ cost overall.
+Several other Bayesian approaches considered there also had difficulty.
+There has been recent progress on scalable algorithms for crossed random effects problems improving both the frequentist and Bayesian approaches.
+For regression problems Ghosh et al.
+(2022a) replace standard equation solving by a backfitting algorithm that has $O(N)$ cost per iteration and they give conditions under which the number of iterations to convergence is $O(1)$ as $N\to\infty$.
+See also Ghandwani et al.
+(2023) for regression with random slopes.
+Papaspiliopoulos et al.
+(2020) use a collapsed Gibbs sampler and then give conditions under which it has linear cost in $N$ for the intercept-only crossed random effects regression model.
+Ghosh & Zhong (2021) do the same after weakening a stringent balance assumption.
+We consider a binary response requiring a generalized linear mixed effects model with crossed random effects that incur the high dimensional integration problem mentioned above.
+There are fewer scalable solutions for this problem.
+There is a frequentist approach by Ghosh et al.
+(2022b) and a Bayesian approach by Papaspiliopoulos et al.
+(2023).
+The all-row-column probit given here is simpler than those two approaches and uses much weaker sampling assumptions.
+Ghosh et al.
+(2022b) developed a penalized quasi-likelihood approach to logistic regression on fixed effects and two crossed random effects.
+That approach had iterations costing $O(N)$ each and empirically the number of iterations was $O(1)$.
+They used estimating equations from Breslow & Clayton (1993) which were based on work by Schall (1991) to maximize the marginal likelihood using a Laplace approximation.
+The quantity being estimated is not exactly the maximum likelihood estimate.
+It is a posterior mode corresponding to a not very informative prior, using some plugged-in weights, a quantity that goes back to Stiratelli et al.
+(1984).
+Penalized quasi-likelihood has a bias that can prevent it from being consistent.
+Even with just the intercept and one random effect, it requires the number of levels of that effect as well as the number of observations at each of those levels to diverge to infinity in order to get a consistent estimate.
+With sample sizes $R=N^{\rho}$ and $C=N^{\kappa}$, for $\rho,\kappa\in(0,1)$ they require $\max(\rho+2\kappa,2\rho+\kappa)<2$ and the observation probability for $(x_{ij},Y_{ij})$ can only vary over a narrow range.
+Papaspiliopoulos et al.
+(2023) extend the collapsed Gibbs sampler to obtain scalable Bayesian inference in generalized linear mixed models with crossed random effects using a reparameterization called local centering.
+They include an intercept and $K\geq 2$ crossed random effects, while also discussing how fixed effects could be incorporated.
+Their approach requires a stringent balance assumption.
+For our data we would need $N_{i\text{\tiny$\bullet$}}=N/C$ for all rows $i=1,\dots,R$ and $N_{\text{\tiny$\bullet$}j}=N/R$ for all $j=1,\dots,C$.
+With this condition, their cost per iteration is $O(N+R+C)$ in our notation.
+The total cost is this cost per iteration times a relaxation time.
+For $K=2$ random effects and a discrete response like we have, they introduce an auxiliary relaxation time $T_{\text{aux}}$ and show in Theorem 1 that the cost is $O[\max(N,R+C)\min\{2N/(R+C),T_{\text{aux}}\}]$.
+Our problem has $\max(R,C)\ll N$ and then their cost is linear in $N$ if and only if $T_{\text{aux}}=O(1)$.
+They get $T_{\text{aux}}=O(1)$ when the observation pattern is uniformly distributed over all patterns with $N_{i\text{\tiny$\bullet$}}=N/C$ and $N_{\text{\tiny$\bullet$}j}=N/R$.
+Both $R$ and $C$ much grow linearly with $N$.
+They then require $N_{i\text{\tiny$\bullet$}}=O(1)$ and $N_{\text{\tiny$\bullet$}j}=O(1)$.
+Our balance criteria are minimal.
+Our main results require $\max(\max_{i}N_{i\text{\tiny$\bullet$}},\max_{j}N_{\text{\tiny$\bullet$}j})/N=o(1)$.
+Both $R$ and $C$ can grow with $N$ at different rates.
+Our simulations, but not our theory, use a balance condition with nearly equal values among the $N_{i\text{\tiny$\bullet$}}$ and nearly equal values among the $N_{\text{\tiny$\bullet$}j}$.
+We study the probit model because its Gaussian latent variable is a good match for Gaussian random effects.
 The probit and logit link functions are nearly proportional outside of tail regions (Agresti, 2002, pp.
 246–247), and so they often give similar results.
-We have some further remarks about the logistic case in our conclusions.
+Crossed models also differ from hierarchical models in that it is only recently that the maximum likelihood estimate for generalized linear mixed models has been shown to be consistent for crossed random effects.
+This was accomplished by the subset argument of Jiang (2013) showing that the score equations have a root that is consistent for the parameter.
+One of the main open questions in generalized linear mixed models, like the ones we consider here, is at what rate do the estimators of the model parameters converge?
+Intuitively one might expect $O(N^{-1/2})$ or $O\{\min(R,C)^{-1/2}\}$, in accordance with the results valid for nested designs (Jiang et al., 2022).
+In some settings estimators for different parameters converge at different rates.
+See Jiang (2013) for some discussion and Lyu et al.
+(2024) for recent work assuming balanced sampling.
+2 The All-Row-Column Method
+2.1 Preliminaries
 We consider two crossed random effects.
-There is a vector $\bm{a}\in\mathbb{R}^{R}$ with elements $a_{i}$ and another vector $\bm{b}\in\mathbb{R}^{C}$ with elements $b_{j}$.
-Conditionally on $\bm{a}$ and $\bm{b}$, the $Y_{ij}$ are independent with
-$\Pr(Y_{ij}=1\!\mid\!\bm{a},\bm{b})=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j}),$ (1)
+There is a vector ${a}\in\mathbb{R}^{R}$ with elements $a_{i}$ and another vector ${b}\in\mathbb{R}^{C}$ with elements $b_{j}$.
+Conditionally on ${a}$ and ${b}$, the $Y_{ij}$ are independent with
+$\mathrm{pr}(Y_{ij}=1\mid{a},{b})=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j}),$ (1)
 where $\Phi(\cdot)$ is the standard normal cumulative distribution function.
-In these models, random effects are typically assumed to be uncorrelated, $\bm{a}\sim\mathcal{N}(0,\sigma^{2}_{A}I_{R})$ independently of $\bm{b}\sim\mathcal{N}(0,\sigma^{2}_{B}I_{C}),$ where $I_{n}$ is the $n\times n$ identity matrix, see, for example, McCullagh and Nelder (1989, page 444).
-The probit model has a representation in terms of latent variables $\varepsilon_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,1)$, via
-$Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j}+\varepsilon_{ij}>0\},$ (2)
-where $1\{E\}$ is the indicator function of the event $E$.
-The probability (1) and the likelihoods derived from it are all conditional on the values of $\bm{x}_{ij}$.
+In these models, random effects are typically assumed to be uncorrelated, ${a}\sim{N}(0,\sigma^{2}_{A}I_{R})$ independently of ${b}\sim{N}(0,\sigma^{2}_{B}I_{C}),$ where $I_{n}$ is the $n\times n$ identity matrix.
+See, for example, McCullagh & Nelder (1989, page 444).
+The probit model has a representation in terms of latent variables $\varepsilon_{ij}\sim{N}(0,1)$, via
+$Y_{ij}=1\{{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j}+\varepsilon_{ij}>0\},$ (2)
+where $1\{E\}$ is the indicator function of the event $E$. The probability (1) and the likelihoods derived from it are all conditional on the values of ${x}_{ij}$.
+We write $\mathcal{S}\subset\{1,\dots,R\}\times\{1,\dots,C\}$ for the set of $(i,j)$ pairs where $({x}_{ij},Y_{ij})$ was observed.
+We also work conditionally on $\mathcal{S}$.
+In our motivating applications the pattern of observation/missingness could be informative.
+Addressing that issue would necessarily require information from outside the data.
+Furthermore, the scaling problem is still the subject of current research even in the non-informative missingness setting.
+Therefore we consider estimation strategies without taking account of missingness.
+The likelihood for ${\theta}=({\beta}^{\mathrm{\scriptscriptstyle T}},\sigma^{2}_{A},\sigma^{2}_{B})^{\mathrm{\scriptscriptstyle T}}$ is a cumbersome integral of size $R+C$,
+$L({\theta})=\sigma_{A}^{-R}\sigma_{B}^{-C}\int_{\mathbb{R}^{R+C}}L({\beta}\mid{a},{b})\prod_{i=1}^{R}\varphi\Bigl{(}\frac{a_{i}}{\sigma_{A}}\Bigr{)}\prod_{j=1}^{C}\varphi\Bigl{(}\frac{b_{j}}{\sigma_{B}}\Bigr{)}\,\mathrm{d}{a}\,\mathrm{d}{b},$ (3)
+where $\varphi(\cdot)$ is the standard normal probability density function, and $L({\beta}\mid{a},{b})$ is the conditional likelihood of ${\beta}$ given the random effects.
+The conditional likelihood we need is
+$L({\beta}\mid{a},{b})=\prod_{(i,j)\in\mathcal{S}}\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j})^{y_{ij}}\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}-a_{i}-b_{j})^{1-y_{ij}},$
+and $L(\theta)$ is commonly called the marginal likelihood.
+The first-order Laplace approximation is a standard approach to approximate the integral in the marginal likelihood.
+The Laplace algorithm maximizes the logarithm of the integrand in the marginal likelihood (3) over ${a}\in\mathbb{R}^{R}$ and ${b}\in\mathbb{R}^{C}$ for fixed ${\theta}$.
+It then multiplies the maximum value of the integrand by $\det\{H^{-1/2}(\theta)\}$ where $H(\theta)\in\mathbb{R}^{(R+C)\times(R+C)}$ is the Hessian of the log integrand with respect to ${a}$ and ${b}$ for fixed ${\theta}$.
+The result is an approximate marginal likelihood $\tilde{L}({\theta})$ that is optimized to get $\hat{{\theta}}$.
+See for example Ogden (2021) or Shun & McCullagh (1995).
+If the square root of the inverse Hessian is computed by standard methods then that alone has cost of $\Omega(N^{3/2})$ by the argument for generalised linear models discussed in §1.
+Similarly, if the inner optimization is done using Newton steps, that will have a cost that is $\Omega(N^{3/2})$ per iteration.
+We return to this issue in §3.2 where a Laplace approximation shows superlinear cost that is $o(N^{3/2})$.
+Even if the Laplace approximation were computable for large $N$, it does not provide asymptotically valid results in our context because the size of the likelihood integral corresponds to the number of random effects $\Omega(N^{1/2})$ and therefore grows too fast with the sample size to ensure consistent results.
+See Shun & McCullagh (1995), Ogden (2021) and Tang & Reid (2024) for a discussion of the conditions that must be satisfied to make the Laplace approximation reliable when the size of the integral grows with the sample size.
+2.2 Scalable composite likelihood inference
+Our approach for a consistent and scalable estimator in high-dimensional probit models with crossed-random effects combines estimates for three misspecified probit models.
+Each of them is constructed through the omission of some random effects.
 By combining (1) and (2) we find that marginally
-$\Pr(Y_{ij}=1)=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})$ (3)
-for $\bm{\gamma}=\bm{\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}$.
-The marginal model (3) can be fit in $\mathcal{O}(N)$ computation, but we then need estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to get the scale right.
-While $\mathrm{sign}(\bm{\gamma}_{k})=\mathrm{sign}(\beta_{k})$ for $k=1,\dots,p$ confidence intervals for $\bm{\gamma}_{k}$ based on model (3) would be naive if they did not account for the dependence among the responses.
-We can similarly find that
-$\Pr(Y_{ij}=1\!\mid\!\bm{a})$ $=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{A}+u_{i}),\quad\text{and}$ (4)
-$\Pr(Y_{ij}=1\!\mid\!\bm{b})$ $=\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{B}+v_{j})$ (5)
-where $\bm{u}\sim\mathcal{N}(0,\tau^{2}_{A}I_{R})$ and $\bm{v}\sim\mathcal{N}(0,\tau^{2}_{B}I_{C})$ for $\tau^{2}_{A}=\sigma^{2}_{A}/(1+\sigma^{2}_{B})$, $\tau^{2}_{B}=\sigma^{2}_{B}/(1+\sigma^{2}_{A})$, $\bm{\gamma}_{A}=\bm{\gamma}(1+\tau^{2}_{A})^{1/2}$ and $\bm{\gamma}_{B}=\bm{\gamma}(1+\tau^{2}_{B})^{1/2}$.
-We fit models (3), (4) and (5) together to estimate our probit model, using the latter two to estimate $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$.
+$\mathrm{pr}(Y_{ij}=1)=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}),$ (4)
+for ${\gamma}={\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}$.
+The proposed method begins with estimation of ${\gamma}$ from the naïve model (4) that omits both of the random effects through maximization of the likelihood
+$L_{\text{all}}({\gamma})=\prod_{(i,j)\in\mathcal{S}}\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})^{y_{ij}}\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})^{1-y_{ij}}.$ (5)
+Maximization of this likelihood requires neither high dimensional integrals nor expensive algebra and we see it taking $O(N)$ computation in our examples.
+We then need estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to get the scale right.
+Our analysis of (4) will also account for within-row and within-column correlations among the $Y_{ij}$.
+While $\mathrm{sign}({\gamma}_{k})=\mathrm{sign}(\beta_{k})$ $(k=1,\dots,p),$ confidence intervals for ${\gamma}_{k}$ based on model (4) would be naïve if they did not account for the dependence among the responses.
+Consider the reparameterization ${\psi}=({\gamma}^{\mathrm{\scriptscriptstyle T}},\tau^{2}_{A},\tau^{2}_{B})^{\mathrm{\scriptscriptstyle T}},$ where
+${\gamma}=\frac{{\beta}}{(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}},\quad\tau^{2}_{A}=\frac{\sigma^{2}_{A}}{1+\sigma^{2}_{B}},\quad\tau^{2}_{B}=\frac{\sigma^{2}_{B}}{1+\sigma^{2}_{A}}.$ (6)
+After dividing ${x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j}+\varepsilon_{ij}$ by $(1+\sigma^{2}_{A})^{1/2}$ or by $(1+\sigma^{2}_{B})^{1/2}$, model (2) implies that
+$\mathrm{pr}(Y_{ij}=1\mid{a})$ $=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{A}+u_{i}),$ (7)
+$\mathrm{pr}(Y_{ij}=1\mid{b})$ $=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{B}+v_{j}),$ (8)
+where ${u}\sim{N}(0,\tau^{2}_{A}I_{R})$ and ${v}\sim{N}(0,\tau^{2}_{B}I_{C})$ for ${\gamma}_{A}={\gamma}(1+\tau^{2}_{A})^{1/2}$ and ${\gamma}_{B}={\gamma}(1+\tau^{2}_{B})^{1/2}$.
+Given $\hat{{\gamma}}$, the maximizer of (5), we proceed with estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the two probit models (7) and (8) that each omit one of the random effects.
 Fitting models (7) and (8) involve simpler integrals than (1) since the latent variable representations of these models
-$Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{A}+u_{i}+\varepsilon_{ij}>0\}\quad\text{and}\quad Y_{ij}=1\{\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma}_{B}+v_{j}+\varepsilon_{ij}>0\}$
-have hierarchical (not crossed) error structures. This is where we are able to replace an $R+C$-dimensional integral by $R+C$ univariate ones.
-We call our method ‘all-row-column’ (ARC).
-The name comes from model (3) that uses all the data at once, model (4) that combines likelihood contribution from within each row and model (5) that combines likelihood contributions from within each column.
+$Y_{ij}=1\{{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{A}+u_{i}+\varepsilon_{ij}>0\},\quad Y_{ij}=1\{{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{B}+v_{j}+\varepsilon_{ij}>0\},$
+have hierarchical (not crossed) error structures.
+This is where we are able to replace the $(R+C)$-dimensional integral (3) by $R+C$ univariate ones.
+Model (7) is fitted by maximizing the row-wise likelihood
+$L_{\text{row}}(\tau^{2}_{A})=\tau_{A}^{-R}\prod_{i=1}^{R}\int_{\mathbb{R}}L_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}_{A}\mid u_{i})\varphi\Bigl{(}\frac{u_{i}}{\tau_{A}}\Bigr{)}\,\mathrm{d}u_{i},$ (9)
+where $L_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}_{A}\mid u_{i})$ is the conditional likelihood of $\hat{{\gamma}}_{A}=\hat{{\gamma}}(1+\tau^{2}_{A})^{1/2}$ given $u_{i},$
+$L_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}_{A}\mid u_{i})=\prod_{j\mid i}\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}_{A}+u_{i})^{y_{ij}}\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}_{A}-u_{i})^{1-y_{ij}},$ (10)
+where $j\mid i={\{j:(i,j)\in\mathcal{S}\}}$ is the set of indices $j$ such that $({x}_{ij},Y_{ij})$ is observed.
+The row-wise likelihood (9) is a product of $R$ one-dimensional integrals and is a function of $\tau^{2}_{A}$ only, because we fix ${\gamma}$ at the estimate $\hat{{\gamma}}$ obtained from the maximization of the all likelihood at the previous step.
+Rows with a single observation do not contribute to estimation of $\tau^{2}_{A}$ because $\mathrm{pr}(Y_{ij}=1)=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}),$ which does not depend on $\tau^{2}_{A}.$ Reversing the roles of the rows and columns, we get a column-wise likelihood $L_{\text{col}}$ which we maximize to get an estimate $\hat{\tau}^{2}_{B}$ of $\tau^{2}_{B}$.
+Finally, we invert the equations in (6) to get
+$\hat{{\beta}}=\hat{{\gamma}}(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2},\quad\hat{\sigma}_{A}^{2}=\frac{\hat{\tau}_{A}^{2}(1+\hat{\tau}^{2}_{B})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}},\quad\hat{\sigma}_{B}^{2}=\frac{\hat{\tau}_{B}^{2}(1+\hat{\tau}^{2}_{A})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}.$
+Note that by the definitions in (6), $\tau^{2}_{A}\tau^{2}_{B}<1$.
+In our computations, we never encountered a setting where $\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}\geq 1$ and so our estimates $\hat{\sigma}^{2}_{A}$ and $\hat{\sigma}^{2}_{B}$ were never negative.
+We call our method ‘all-row-column’.
+The name comes from model (4) that uses all the data at once, model (7) that combines likelihood contributions from within each row and model (8) that combines likelihood contributions from within each column.
+Figure 1 illustrates those models for a data set of $N=39$ observations in $R=10$ rows and $C=10$ columns.
+In each of three misspecified probit models, points in different boxes are assumed to be independent.
+The left panel of Figure 1 illustrates the ‘all’ model with independent data.
+The next two illustrate the ‘row’ and ‘column’ hierarchical models.
+We combine fits from these three misspecified models to get consistent scalable formulas despite the dependencies involved.
 This approach is a new form of composite likelihood (Lindsay, 1988; Varin et al., 2011).
-An earlier version of composite likelihood that was applied to crossed random effects is discussed in Bellio and Varin (2005).
+An earlier version of composite likelihood that was applied to crossed random effects is discussed in Bellio & Varin (2005).
 They considered a standard pairwise likelihood which is however not scalable and thus inappropriate for our problem.
-Instead, our ARC method does not only consistently estimates $\bm{\beta}$, $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ but it is also scalable having $\mathcal{O}(N)$ cost.
-An outline of this paper is as follows: Section 2 develops further notation and properties for the models we combine.
-Section 3 presents our ARC method and develops a robust sandwich estimate of the variance of the regression parameter estimates.
-Section 4 presents some simulations of the ARC method.
-We also compare it to an infeasible oracle method that estimates $\bm{\gamma}$ from (3) and then adjusts the estimate using the true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$.
-A second comparison is to the most popular frequentist method for fitting generalized linear models with random effects, that is the first-order Laplace approximate maximum likelihood estimator described in Skaug (2002) and implemented in the popular R (R Core Team, 2024) package lme4 (Bates et al., 2015).
-Section 5 illustrates our proposed method on real electronic commerce data provided by Stitch Fix.
-Section 6 gives some bibliographic references focussing on the challenges that the crossed random effects setting brings.
-Section 7 has our conclusions.
-Finally, there is a supplement with proofs of consistent estimation, more references, consideration of pairwise composite likelihood and a more detailed discussion of the simulation results.
-2 Probit Regression with Crossed Random Effects
-As described in the introduction we have responses $Y_{ij}\in\{0,1\}$, covariates $\bm{x}_{ij}\in\mathbb{R}^{p}$ and random effects $\bm{a}\sim\mathcal{N}(0,\sigma^{2}_{A}I_{R})$, $\bm{b}\sim\mathcal{N}(0,\sigma^{2}_{B}I_{C})$.
-The random effects are independent of each other, and we condition on the observed values of the covariates $\bm{x}_{ij}$, so we then treat them as non-random.
-We write $\mathcal{S}\subset\{1,\dots,R\}\times\{1,\dots,C\}$ for the set of $(i,j)$ pairs where $(\bm{x}_{ij},Y_{ij})$ was observed.
-We also work conditionally on $\mathcal{S}$.
-In our motivating applications the pattern of observation/missingness could well be informative.
-Addressing that issue would necessarily require information from outside the data.
-Furthermore, the scaling problem is still the subject of current research in the non-informative missingness setting.
-Therefore we consider estimation strategies without taking account of missingness.
-We choose to study the probit model (1) where, conditionally on $\bm{a}$ and $\bm{b}$, $Y_{ij}\in\{0,1\}$ are independent with $\Pr(Y_{ij}=1\!\mid\!\bm{a},\bm{b})=\Phi(\bm{x}_{ij}^{\mathsf{T}}\beta+a_{i}+b_{j})$.
-The full likelihood for $\bm{\theta}=(\bm{\beta}^{\mathsf{T}},\sigma^{2}_{A},\sigma^{2}_{B})^{\mathsf{T}}$ is the integral
-$L(\bm{\theta})=\sigma_{A}^{-R}\sigma_{B}^{-C}\int_{\mathbb{R}^{R+C}}L(\bm{\beta}\!\mid\!\bm{a},\bm{b})\prod_{i=1}^{R}\varphi\Bigl{(}\frac{a_{i}}{\sigma_{A}}\Bigr{)}\prod_{j=1}^{C}\varphi\Bigl{(}\frac{b_{j}}{\sigma_{B}}\Bigr{)}\,\mathrm{d}\bm{a}\,\mathrm{d}\bm{b},$ (6)
-where $\varphi(\cdot)$ is the standard normal probability density function, and $L(\bm{\beta}\!\mid\!\bm{a},\bm{b})$ is the conditional likelihood of $\bm{\beta}$ given the random effects.
-The conditional likelihood we need is
-$L(\bm{\beta}\!\mid\!\bm{a},\bm{b})=\prod_{(i,j)\in\mathcal{S}}\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}+a_{i}+b_{j})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\beta}-a_{i}-b_{j})^{1-y_{ij}}.$
-An option to alleviate the computational cost of likelihood computations is provided by composite likelihood methods (Lindsay, 1988).
-Composite likelihoods are inference functions constructed by multiplying component likelihoods.
-See Varin et al.
-(2011) for a review.
-Several articles discussed composite likelihood estimation of models for dependent binary data; see, for example, Heagerty and Lele (1998), Cox and Reid (2004) or Kuk (2007).
-All these works consider a particular type of composite likelihood called pairwise likelihood that is built taking the product of bivariate marginal probabilities.
-Bellio and Varin (2005) investigated pairwise likelihood inference in logistic regression with crossed random effects.
-Their pairwise likelihood involves all the pairs of correlated observations, that is those pairs that share the row- or the column-random effect.
-Pairwise likelihood is not attractive in this setting.
-The number of pairs of data points far exceeds the $\mathcal{O}(N)$ constraint.
-In a setting with $R$ rows, the average row has $N/R$ elements in it.
-If all the rows had that many elements then the number of pairs would be $\Omega(R(N/R)^{2})=\Omega(N^{2}/R)$.
-Unequal numbers of observations per row can only increase this count.
-As a result the cost must be $\Omega(\max(N^{2}/R,N^{2}/C))$.
-In our motivating settings, both $R$ and $C$ are $o(N)$ so the cost cannot then be $\mathcal{O}(N)$.
-A possible remedy to the lack of scalability would be to use only $\mathcal{O}(N)$ randomly sampled pairs of correlated observations.
-Although sampling might be a valid option, we decided to avoid this approach.
-First, we prefer estimates that do not depend on unnecessary randomness in the algorithm.
-Second, some preliminary simulations reported in Section S2 of the Supplementary Material indicate that the composite likelihood discussed in the next section satisfy the $\mathcal{O}(N)$ constraint with a statistical efficiency comparable to that of the pairwise likelihood without sampling of the pairs, which is however not feasible to compute in the high-dimensional applications that motivate this work.
-Third, we view our ARC method as new form of composite likelihood using all pairs, all triples and indeed all subsets of observations within each of the rows, and also within each of the columns.
-It thus captures all the data dependencies already included in pairwise likelihood and more while only requiring three estimating equations and $\mathcal{O}(N)$ cost.
-3 The All-Row-Column Method
-3.1 Three misspecified probit models
-Our ARC approach for scalable inference in high-dimensional probit models with crossed-random effects combines estimates for three misspecified probit models.
-Each of them is constructed through omission of some random effects.
-Consider the reparameterization $\bm{\psi}=(\bm{\gamma}^{\mathsf{T}},\tau^{2}_{A},\tau^{2}_{B})^{\mathsf{T}},$ whose components were already anticipated in Section 1, namely,
-$\bm{\gamma}=\frac{\bm{\beta}}{(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}},\quad\tau^{2}_{A}=\frac{\sigma^{2}_{A}}{1+\sigma^{2}_{B}}\quad\text{and}\quad\tau^{2}_{B}=\frac{\sigma^{2}_{B}}{1+\sigma^{2}_{A}}.$ (7)
-The proposed method begins with estimation of $\bm{\gamma}$ from the naïve model (3) that omits both of the random effects through maximization of the likelihood of the probit model
-$L_{\text{all}}(\bm{\gamma})=\prod_{(i,j)\in\mathcal{S}}\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})^{1-y_{ij}}.$ (8)
-Given $\hat{\bm{\gamma}}$, the maximizer of (8), we proceed with estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the two probit models (4) and (5) that each omit one of the random effects.
-Accordingly, parameter $\tau^{2}_{A}$ is estimated from the row-wise likelihood
-$L_{\text{row}}(\tau^{2}_{A})=\tau_{A}^{-R}\prod_{i=1}^{R}\int_{\mathbb{R}}L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})\varphi\Bigl{(}\frac{u_{i}}{\tau_{A}}\Bigr{)}\,\mathrm{d}u_{i},$ (9)
-where $L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})$ is the conditional likelihood of $\hat{\bm{\gamma}}_{A}=\hat{\bm{\gamma}}(1+\tau^{2}_{A})^{1/2}$ given $u_{i},$
-$L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})=\prod_{j|i}\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{A}+u_{i})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{A}-u_{i})^{1-y_{ij}},$ (10)
-where $j|i={\{j:(i,j)\in\mathcal{S}\}}$ is the set of indices $j$ such that $(\bm{x}_{ij},Y_{ij})$ is observed.
-Notice that the row-wise likelihood is a function of $\tau^{2}_{A}$ only, because we fix $\bm{\gamma}$ at the estimate $\hat{\bm{\gamma}}$ obtained from the maximization of the all likelihood at the previous step.
-Rows with a single observation do not contribute to estimation of $\tau^{2}_{A}$ because $\Pr(Y_{ij}=1)=\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}),$ which does not depend on $\tau^{2}_{A}.$
-Computation of the row-wise likelihood requires one to approximate $R$ univariate integrals of a standard hierarchical probit model.
-Since $R$ is large in the applications that motivate this work, then accurate approximation of the one-dimensional integrals is crucial.
-Otherwise, the accumulation of approximation errors could induce serious biases in the estimation of $\tau^{2}_{A}$.
-Our intensive numerical studies indicated that accurate approximation of the integrals is obtained by the well-established adaptive Gaussian quadrature (Liu and Pierce, 1994) with a suitable choice of the number of quadrature nodes.
-Following Jin and Andersson (2020) and Stringer and Bilodeau (2022), we set the number of quadrature nodes equal to $k(R)=\lceil(3/2)\log_{2}(R)-2\rceil,$ a number rather larger than is typically considered in the GLMM literature.
-For example, in the Stitch Fix application discussed in Section 5, we have $R=741,221$ and therefore we have used $k=28$ quadrature nodes.
-Reversing rows and columns, $\tau^{2}_{B}$ is estimated from the column-wise likelihood
-$L_{\text{col}}(\tau^{2}_{B})=\tau_{B}^{-C}\prod_{j=1}^{C}\int_{\mathbb{R}}L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})\varphi\Bigl{(}\frac{v_{j}}{\tau_{B}}\Bigr{)}\,\mathrm{d}v_{j},$ (11)
-where $L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})$ is the conditional likelihood of $\hat{\bm{\gamma}}_{B}=\hat{\bm{\gamma}}(1+\tau^{2}_{B})^{1/2}$ given $v_{j},$
-$L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})=\prod_{i|j}\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{B}+v_{j})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{B}-v_{j})^{1-y_{ij}}$
-with $i|j={\{i:(i,j)\in\mathcal{S}\}}.$ Finally, the estimates under the original parameterization are obtained by inverting the equations in (7),
-$\hat{\bm{\beta}}=\hat{\bm{\gamma}}(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2},\quad\hat{\sigma}_{A}^{2}=\frac{\hat{\tau}_{A}^{2}(1+\hat{\tau}^{2}_{B})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}\quad\text{and}\quad\hat{\sigma}_{B}^{2}=\frac{\hat{\tau}_{B}^{2}(1+\hat{\tau}^{2}_{A})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}.$
-Note that by the definitions in (7), $\tau^{2}_{A}\tau^{2}_{B}<1$.
-In our computations, we never encountered a setting where $\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}>1$ [should be $\geq$ not $>$] and so our estimates $\hat{\sigma}^{2}_{A}$ and $\hat{\sigma}^{2}_{B}$ were never negative.
-In case of a violation, we suggest taking $\hat{\sigma}^{2}_{A}=\hat{\sigma}^{2}_{B}=0$ as is commonly done for negative variance component estimates in random effect settings.
-The ARC method offers a significant computational saving compared to ordinary maximum likelihood estimation for two reasons.
-In addition to the simplified integration, the reparameterization used by the ARC method conveniently separates estimation of the parameters: regression parameters are estimated from a standard probit model, whereas estimation of the dependence parameters require two distinct univariate optimizations of the row- and column-wise likelihoods.
-In the next Theorem, we prove the weak consistency of the maximizer $\hat{\bm{\gamma}}$ of the ‘all’ likelihood (8) to the value $\bm{\gamma}$ of equation (7).
+Instead, our all-row-column method has $O(N)$ cost per iteration and we get $O(N)$ cost empirically.
+Figure 1: The all-row-column method for $N=39$ observations (marked $\bullet$) in $R=10$ rows and $C=10$ columns.
+Computation of the row- and column-wise likelihoods require one to approximate up to $R$ and $C$ univariate integrals of a standard hierarchical probit model.
+Since $R$ and $C$ are large in the applications that motivate this work, then accurate approximation of the one-dimensional integrals is crucial.
+Otherwise, the accumulation of approximation errors could induce serious biases in the estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$.
+Our intensive numerical studies indicated that accurate approximation of the integrals is obtained by the well-established adaptive Gauss-Hermite quadrature (e.g., Liu & Pierce, 1994) with a suitable choice of the number of quadrature nodes.
+In the next Theorem, we prove the weak consistency of the maximizer $\hat{{\gamma}}$ of the ‘all’ likelihood (5) to the value ${\gamma}$ of equation (6).
 First we introduce some notation.
-Let $Z_{ij}=1$ if $(\bm{x}_{ij},Y_{ij})$ is observed and take $Z_{ij}=0$ otherwise.
+Let $Z_{ij}=1$ if $({x}_{ij},Y_{ij})$ is observed and take $Z_{ij}=0$ otherwise.
 The number of observations in row $i$ is $N_{i\text{\tiny$\bullet$}}=\sum_{j=1}^{C}Z_{ij}$ and similarly column $j$ has $N_{\text{\tiny$\bullet$}j}=\sum_{i=1}^{R}Z_{ij}$ observations in it.
 Let $\epsilon_{R}=\max_{i}N_{i\text{\tiny$\bullet$}}/N$ and $\epsilon_{C}=\max_{j}N_{\text{\tiny$\bullet$}j}/N$.
-In our motivating settings it is reasonable to assume that $\max(\epsilon_{R},\epsilon_{C})\to 0$ as $N\to\infty$.
-When we need a double sum over all pairs of data we use index $r$ as a ‘second $i$’ and $s$ as a ‘second $j$’.
+We assume that $\max(\epsilon_{R},\epsilon_{C})\to 0$ as $N\to\infty$.
 We take $Z_{ij}$ to be deterministic with at most one observation for any $(i,j)$ pair.
-In our motivating applications that would either never happen or be very rare.
-If rare, one could use the most recent observation only.
-The $\bm{x}_{ij}$ are not dependent on the $Z_{ij}$.
-Finally, $Y_{ij}$ are sampled from their probit distribution conditionally on $\bm{x}_{ij}$.
-Theorem 1.
-Let $Y_{ij}\in\{0,1\}$ follow the crossed random effects probit model (1) with true value $\bm{\gamma}_{0}$ for the parameter $\bm{\gamma}=\bm{\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ Let the number of observations $N\to\infty$ while $\max(\epsilon_{R},\epsilon_{C})\to 0$.
-Let $\bm{x}_{ij}\in\mathbb{R}^{p}$ satisfy
+In our motivating applications, we would seldom if ever have multiple observations for one $(i,j)$ pair.
+Even then, one might only use the most recent of those observations.
+The ${x}_{ij}$ are not dependent on the $Z_{ij}$.
+Finally, $Y_{ij}$ are sampled from their probit distribution conditionally on ${x}_{ij}$.
+Theorem 1
+Let $Y_{ij}\in\{0,1\}$ follow the crossed random effects probit model (1) with true value ${\gamma}_{0}$ for the parameter ${\gamma}={\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ Let the number of observations $N\to\infty$ while $\max(\epsilon_{R},\epsilon_{C})\to 0$.
+Let ${x}_{ij}\in\mathbb{R}^{p}$ satisfy
   1. 1.
-    $\|\bm{x}_{ij}\|\leq B<\infty;$
+    $\|{x}_{ij}\|\leq B<\infty;$
   2. 2.
-    $N^{-1}\sum_{ij}Z_{ij}\bm{x}_{ij}\bm{x}_{ij}^{\mathsf{T}}\to V\in\mathbb{R}^{p\times p},$ where $V$ is positive definite;
+    $N^{-1}\sum_{ij}Z_{ij}{x}_{ij}{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\to V\in\mathbb{R}^{p\times p},$ where $V$ is positive definite;
   3. 3.
-    there is no nonzero vector $\bm{v}\in\mathbb{R}^{p}$ such that $\bm{v}^{\mathsf{T}}\bm{x}_{ij}\geq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=1$ and $\bm{v}^{\mathsf{T}}\bm{x}_{ij}\leq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=0$.
-Let $\hat{\bm{\gamma}}\in\mathbb{R}^{p}$ be any maximizer of (8). Then for any $\epsilon>0$
-$\Pr(\|\hat{\bm{\gamma}}-\bm{\gamma}_{0}\|>\epsilon)\to 0,\;\text{as $N\to\infty$}.$
-The proof is reported in Section S1 of the Supplementary Material and we obtained it by adapting the proof strategy of Lumley and Mayer Hamblett (2003) to our setting.
+    there is no nonzero vector ${v}\in\mathbb{R}^{p}$ such that ${v}^{\mathrm{\scriptscriptstyle T}}{x}_{ij}\geq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=1$ and ${v}^{\mathrm{\scriptscriptstyle T}}{x}_{ij}\leq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=0$.
+Let $\hat{{\gamma}}\in\mathbb{R}^{p}$ be any maximizer of (5). Then for any $\epsilon>0$
+$\mathrm{pr}(\|\hat{{\gamma}}-{\gamma}_{0}\|>\epsilon)\to 0,\;\text{as $N\to\infty$}.$
+The proof is reported in Section S1 of the Supplementary Material and we obtained it by adapting the proof strategy of Lumley & Mayer Hamblett (2003) to our setting.
+The balance condition that no single row or column have a non-vanishing fraction of data is much weaker than the norm for the crossed random effects literature.
 Now we consider consistent estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the row and column likelihoods, respectively.
-Those provide consistent estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ with which one can then adjust the consistent estimate of $\bm{\gamma}$ to get a consistent estimate of $\bm{\beta}$.
-Theorem 2.
+Those provide consistent estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ with which one can then adjust the consistent estimate of ${\gamma}$ to get a consistent estimate of ${\beta}$.
+Theorem 2
 Under the assumptions of Theorem 1, there is a root of the row likelihood equation that is a consistent estimator for $\tau^{2}_{A}$ and a root of the column likelihood equation that is a consistent estimator for $\tau^{2}_{B}.$
-Our proof of this theorem is also reported in Section S1 of the Supplementary Material.
-The proof uses the subset argument of Jiang (2013) to show Cramer consistency of the MLE row likelihood and equivalently of the column likelihood.
-If one does not find Cramer consistency sufficient, then it is possible to construct a consistent estimator using all of the data from a subset of rows with a large number of observations.
-In Section S1 of the Supplementary Material we show how to construct such a consistent estimator and we give conditions under which the number of large rows will diverge to infinity as $N\to\infty$.
-See Theorem 3 of the Supplementary Material.
-We prefer our ARC estimator to an approach using just large rows, because it would be awkward to have to decide in practice which rows to use, and because we believe that there is valuable information in the other, smaller, rows.
+Our proof of this theorem is also reported in §S1 of the Supplementary Material.
+The proof uses the subset argument of Jiang (2013) to show Cramer consistency of the maximum row likelihood estimator of $\tau^{2}_{A}$ and equivalently of the maximum column likelihood estimator of $\tau^{2}_{B}$.
 2.3 Robust sandwich variance
-After a customary Taylor approximation, the variance of $\hat{\bm{\theta}}$ is $\mathrm{var}(\hat{\bm{\theta}})\doteq D\,\mathrm{var}(\hat{\bm{\psi}})\,D^{\mathsf{T}},$ with $D$ the Jacobian matrix of the reparameterization from $\bm{\theta}$ to $\bm{\psi}$.
-Letting $\bm{u}_{\text{arc}}(\bm{\psi})$ denote the score vector of the ARC estimator constructed by stacking the scores of the three misspecified likelihoods, the asymptotic variance of $\hat{\bm{\psi}}$ is
-$\mathrm{avar}(\hat{\bm{\psi}})=\mathcal{I}_{\text{arc}}^{-1}(\bm{\psi}_{0})V_{\text{arc}}(\bm{\psi}_{0})\mathcal{I}_{\text{arc}}^{-1}(\bm{\psi}_{0}),$ (12)
-where $\mathcal{I}_{\text{arc}}({\bm{\psi}})=-\mathrm{E}(\partial\bm{u}_{\text{arc}}({\bm{\psi}})/\partial{\bm{\psi}})$ and $V_{\text{arc}}({\bm{\psi}})=\mathrm{var}(\bm{u}_{\text{arc}}({\bm{\psi}}))$ are the expected information and the score variance for the ARC estimator.
-These two matrices are not equal because the second Bartlett identity does not hold for misspecified likelihoods that constitute the ARC method.
-While estimation of the ‘bread’ matrix $\mathcal{I}_{\text{arc}}$ of the sandwich is not problematic, direct computation of the ‘filling‘ matrix $V_{\text{arc}}$ is not feasible in our large-scale setup because it requires us to approximate a large number of multidimensional integrals with a cost that does not meet our $\mathcal{O}(N)$ constraint.
-Since ARC estimates require $\mathcal{O}(N)$ computations, we can estimate the variance of $\hat{\bm{\theta}}$ with a parametric bootstrap.
+After a customary Taylor approximation, the variance of $\hat{{\theta}}$ is $\mathrm{var}(\hat{{\theta}})\doteq D\,\mathrm{var}(\hat{{\psi}})\,D^{\mathrm{\scriptscriptstyle T}},$ with $D$ the Jacobian matrix of the reparameterization from ${\theta}$ to ${\psi}$.
+Letting $\psi_{0}$ denote the true value of $\psi$ and ${u}_{\text{arc}}({\psi})$ denote the score vector of the all-row-column estimator constructed by stacking the scores of the three misspecified likelihoods, the asymptotic variance of $\hat{{\psi}}$ is
+$\mathrm{avar}(\hat{{\psi}})=\mathcal{I}_{\text{arc}}^{-1}({\psi}_{0})V_{\text{arc}}({\psi}_{0})\mathcal{I}_{\text{arc}}^{-1}({\psi}_{0}),$ (11)
+where $\mathcal{I}_{\text{arc}}(\psi)=-\mathrm{E}\{\partial{u}_{\text{arc}}(\psi)/\partial\psi\}$ and $V_{\text{arc}}(\psi)=\mathrm{var}\{u_{\text{arc}}(\psi)\}$ are the expected information and the score variance for the all-row-column estimator.
+These two matrices are not equal because the second Bartlett identity does not hold for the misspecified likelihoods that constitute the all-row-column method.
+While estimation of the ‘bread’ matrix $\mathcal{I}_{\text{arc}}$ of the sandwich is not problematic, direct computation of the ‘filling’ matrix $V_{\text{arc}}$ is not feasible in our large-scale setup because it requires us to approximate a large number of multidimensional integrals with a cost that does not meet our $O(N)$ constraint.
+Since all-row-column estimates require $O(N)$ computations, we can estimate the variance of $\hat{\theta}$ with a parametric bootstrap.
 However, it is preferable to evaluate the estimation uncertainty without assuming the correctness of the fitted model, for example using the nonparametric pigeonhole bootstrap described in Owen (2007).
@@
 It does not require homoscedasticity of either the random effects or the errors.
-Now we describe a convenient approach that we have developed to estimate the variance of $\hat{\bm{\beta}}$ in $\mathcal{O}(N)$ operations without the need of resampling and repeated fitting as in the bootstraps mentioned above.
-The partitioned expected ARC information matrix is
-$\mathcal{I}_{\text{arc}}(\bm{\psi})$ $=-\begin{pmatrix}\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial\bm{\gamma}\partial\bm{\gamma}^{\mathsf{T}})&0&0\\ \mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\bm{\gamma}\partial\tau^{2}_{A})&\mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\tau^{2}_{A}\partial\tau^{2}_{A})&0\\ \mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\bm{\gamma}\partial\tau^{2}_{B})&0&\mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\tau^{2}_{B}\partial\tau^{2}_{B})\\ \end{pmatrix}.$
-Since the expected ARC information is triangular, then the asymptotic variance for $\hat{\bm{\gamma}}$ is
-$\mathrm{avar}(\hat{\bm{\gamma}})=\mathcal{I}_{\text{all}}^{-1}(\bm{\gamma}_{0})V_{\text{all}}(\bm{\gamma}_{0})\mathcal{I}_{\text{all}}^{-1}(\bm{\gamma}_{0}),$ (13)
-where $\mathcal{I}_{\text{all}}({\bm{\gamma}})=-\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial\bm{\gamma}\partial\bm{\gamma}^{\mathsf{T}})$ and $V_{\text{all}}({\bm{\gamma}})=\mathrm{var}(\partial\ell_{\text{all}}/\partial\bm{\gamma})$ are the Fisher expected information and the score variance of the marginal probit model.
-The asymptotic variance (13) for $\hat{\bm{\gamma}}$ is thus the same as that of the estimator that maximizes the all likelihood when the nuisance parameters $\tau^{2}_{A}$ and $\tau^{2}_{B}$ are known.
-The robust sandwich estimator of the variance of $\hat{\bm{\gamma}}$ is obtained by replacing $\mathcal{I}_{\text{all}}(\bm{\gamma}_{0})$ and $V_{\text{all}}(\bm{\gamma}_{0})$ with some estimators that are consistent and robust to misspecification.
+Now we describe a convenient approach that we have developed to estimate the variance of $\hat{{\beta}}$ in $O(N)$ operations without the need of resampling and repeated fitting as in the bootstraps mentioned above.
+The partitioned expected all-row-column information matrix is
+$\mathcal{I}_{\text{arc}}({\psi})$ $=-\begin{pmatrix}\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial{\gamma}\partial{\gamma}^{\mathrm{\scriptscriptstyle T}})&0&0\\[5.0pt] \mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial{\gamma}\partial\tau^{2}_{A})&\mathrm{E}(\partial^{2}\ell_{\text{row}}/\partial\tau^{2}_{A}\partial\tau^{2}_{A})&0\\[5.0pt] \mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial{\gamma}\partial\tau^{2}_{B})&0&\mathrm{E}(\partial^{2}\ell_{\text{col}}/\partial\tau^{2}_{B}\partial\tau^{2}_{B})\\ \end{pmatrix},$
+where $\ell_{\text{all}}=\log L_{\text{all}}(\gamma)$ from (5), $\ell_{\text{row}}=\log L_{\text{row}}(\tau^{2}_{A})$ from (9) and $\ell_{\text{col}}=\log L_{\text{col}}(\tau^{2}_{B})$.
+Since $\mathcal{I}_{\text{arc}}({\psi})$ is triangular, then the asymptotic variance for $\hat{{\gamma}}$ is
+$\mathrm{avar}(\hat{{\gamma}})=\mathcal{I}_{\text{all}}^{-1}({\gamma}_{0})V_{\text{all}}({\gamma}_{0})\mathcal{I}_{\text{all}}^{-1}({\gamma}_{0}),$ (12)
+where $\mathcal{I}_{\text{all}}({{\gamma}})=-\mathrm{E}(\partial^{2}\ell_{\text{all}}/\partial{\gamma}\partial{\gamma}^{\mathrm{\scriptscriptstyle T}})$ and $V_{\text{all}}({{\gamma}})=\mathrm{var}(\partial\ell_{\text{all}}/\partial{\gamma})$ are the Fisher expected information and the score variance of the all likelihood.
+The asymptotic variance (12) for $\hat{{\gamma}}$ is thus the same as that of the estimator that maximizes the all likelihood when the nuisance parameters $\tau^{2}_{A}$ and $\tau^{2}_{B}$ are known.
+The robust sandwich estimator of the variance of $\hat{{\gamma}}$ is obtained by replacing $\mathcal{I}_{\text{all}}({\gamma}_{0})$ and $V_{\text{all}}({\gamma}_{0})$ with some estimators that are consistent and robust to misspecification.
 The expected information is naturally estimated with the observed information,
-$J_{\text{all}}(\hat{\bm{\gamma}})=\sum_{ij}Z_{ij}\varphi(\hat{\eta}_{ij})\left(y_{ij}\frac{\varphi(\hat{\eta}_{ij})+\hat{\eta}_{ij}\,\Phi(\hat{\eta}_{ij})}{\Phi(\hat{\eta}_{ij})^{2}}+(1-y_{ij})\frac{\varphi(\hat{\eta}_{ij})-\hat{\eta}_{ij}\,\Phi(-\hat{\eta}_{ij})}{\Phi(-\hat{\eta}_{ij})^{2}}\right)\bm{x}_{ij}\bm{x}_{ij}^{\mathsf{T}},$
-where $\hat{\eta}_{ij}=\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}.$ Estimation of $V_{\text{all}}(\bm{\gamma}_{0})$ is more involved. This matrix can be decomposed into the sum of three terms,
-$V_{\text{all}}(\bm{\gamma}_{0})$ $=\sum_{ij}\sum_{rs}Z_{ij}Z_{rs}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{rs}^{\mathsf{T}}(\bm{\gamma}_{0})\right)$
-$=\sum_{ijs}Z_{ij}Z_{is}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{is}^{\mathsf{T}}(\bm{\gamma}_{0})\right)+\sum_{ijr}Z_{ij}Z_{rj}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{rj}^{\mathsf{T}}(\bm{\gamma}_{0})\right)$
-$-\sum_{ij}Z_{ij}\mathrm{E}\left(\bm{u}_{ij}(\bm{\gamma}_{0})\bm{u}_{ij}^{\mathsf{T}}(\bm{\gamma}_{0})\right),$
-where $\bm{u}_{ij}(\bm{\gamma})$ is the score for a single observation $Y_{ij},$
-$\bm{u}_{ij}(\bm{\gamma})=\frac{\varphi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\left(y_{ij}-\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\right)\bm{x}_{ij}}{\Phi(\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})\Phi(-\bm{x}_{ij}^{\mathsf{T}}\bm{\gamma})}.$
-The corresponding estimator of $V_{\text{all}}(\bm{\gamma}_{0})$ is
-$\widehat{V}_{\text{all}}(\hat{\bm{\gamma}})=\widehat{V}_{A}(\hat{\bm{\gamma}})+\widehat{V}_{B}(\hat{\bm{\gamma}})-\widehat{V}_{A\cap B}(\hat{\bm{\gamma}}),$
+$J_{\text{all}}(\hat{{\gamma}})=\sum_{ij}Z_{ij}\varphi(\hat{\eta}_{ij})\biggl{\{}y_{ij}\frac{\varphi(\hat{\eta}_{ij})+\hat{\eta}_{ij}\,\Phi(\hat{\eta}_{ij})}{\Phi(\hat{\eta}_{ij})^{2}}+(1-y_{ij})\frac{\varphi(\hat{\eta}_{ij})-\hat{\eta}_{ij}\,\Phi(-\hat{\eta}_{ij})}{\Phi(-\hat{\eta}_{ij})^{2}}\biggr{\}}{x}_{ij}{x}_{ij}^{\mathrm{\scriptscriptstyle T}},$
+where $\hat{\eta}_{ij}={x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}.$ Estimation of $V_{\text{all}}({\gamma}_{0})$ is more involved.
+This matrix can be decomposed into the sum of three terms,
+$V_{\text{all}}({\gamma}_{0})$ $=\sum_{ij}\sum_{rs}Z_{ij}Z_{rs}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{rs}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}}$
+$=\sum_{ijs}Z_{ij}Z_{is}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{is}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}}+\sum_{ijr}Z_{ij}Z_{rj}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{rj}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}}$
+$-\sum_{ij}Z_{ij}\mathrm{E}\bigl{\{}{u}_{ij}({\gamma}_{0}){u}_{ij}^{\mathrm{\scriptscriptstyle T}}({\gamma}_{0})\bigr{\}},$
+where ${u}_{ij}({\gamma})$ is the score for a single observation $Y_{ij},$
+${u}_{ij}({\gamma})=\frac{\varphi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})\bigl{\{}y_{ij}-\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})\bigr{\}}{x}_{ij}}{\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})}.$
+The corresponding estimator of $V_{\text{all}}({\gamma}_{0})$ is
+$\hat{V}_{\text{all}}(\hat{{\gamma}})=\hat{V}_{A}(\hat{{\gamma}})+\hat{V}_{B}(\hat{{\gamma}})-\hat{V}_{A\cap B}(\hat{{\gamma}}),$
 whose components are computed grouping the individual scores with respect to each random effect and their interaction,
-$\widehat{V}_{A}(\hat{\bm{\gamma}})=\sum_{i}\bm{u}_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}})\bm{u}_{i\text{\tiny$\bullet$}}^{\mathsf{T}}(\hat{\bm{\gamma}}),\quad\widehat{V}_{B}(\hat{\bm{\gamma}})=\sum_{j}\bm{u}_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}})\bm{u}_{\text{\tiny$\bullet$}j}^{\mathsf{T}}(\hat{\bm{\gamma}}),\quad\widehat{V}_{A\cap B}(\hat{\bm{\gamma}})=\sum_{ij}Z_{ij}\bm{u}_{ij}(\hat{\bm{\gamma}})\bm{u}_{ij}^{\mathsf{T}}(\hat{\bm{\gamma}}),$
-where $\bm{u}_{i\text{\tiny$\bullet$}}(\bm{\gamma})=\sum_{j}Z_{ij}\bm{u}_{ij}(\bm{\gamma})$ and $\bm{u}_{\text{\tiny$\bullet$}j}(\bm{\gamma})=\sum_{i}Z_{ij}\bm{u}_{ij}(\bm{\gamma}).$ Estimators of the form $\widehat{V}_{\text{all}}(\hat{\bm{\gamma}})$ are used in statistical modeling of data clustered within multiple levels in medical applications (Miglioretti and Heagerty, 2004) and in economics (Cameron et al., 2011) where they are known as two-way cluster-robust sandwich estimators.
-Finally, we approximate the variance of $\hat{\bm{\beta}}$ by plugging in the estimates $\widehat{\sigma}_{A}^{2}$ and $\widehat{\sigma}_{B}^{2}$,
-$\widehat{\mathrm{var}}(\hat{\bm{\beta}})=(1+\widehat{\sigma}_{A}^{2}+\widehat{\sigma}_{B}^{2})\,\widehat{\mathrm{var}}(\hat{\bm{\gamma}})=(1+\widehat{\sigma}_{A}^{2}+\widehat{\sigma}_{B}^{2})\,{J}_{\text{all}}^{-1}(\hat{\bm{\gamma}})\widehat{V}_{\text{all}}(\hat{\bm{\gamma}}){J}_{\text{all}}^{-1}(\hat{\bm{\gamma}}).$ (14)
+$\hat{V}_{A}(\hat{{\gamma}})=\sum_{i}{u}_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}){u}_{i\text{\tiny$\bullet$}}^{\mathrm{\scriptscriptstyle T}}(\hat{{\gamma}}),\quad\hat{V}_{B}(\hat{{\gamma}})=\sum_{j}{u}_{\text{\tiny$\bullet$}j}(\hat{{\gamma}}){u}_{\text{\tiny$\bullet$}j}^{\mathrm{\scriptscriptstyle T}}(\hat{{\gamma}}),\quad\hat{V}_{A\cap B}(\hat{{\gamma}})=\sum_{ij}Z_{ij}{u}_{ij}(\hat{{\gamma}}){u}_{ij}^{\mathrm{\scriptscriptstyle T}}(\hat{{\gamma}}),$
+where ${u}_{i\text{\tiny$\bullet$}}({\gamma})=\sum_{j}Z_{ij}{u}_{ij}({\gamma})$ and ${u}_{\text{\tiny$\bullet$}j}({\gamma})=\sum_{i}Z_{ij}{u}_{ij}({\gamma}).$ Estimators of the form $\hat{V}_{\text{all}}(\hat{{\gamma}})$ are used in statistical modeling of data clustered within multiple levels in medical applications (Miglioretti & Heagerty, 2004) and in economics (Cameron et al., 2011) where they are known as two-way cluster-robust sandwich estimators.
+Finally, we approximate the variance of $\hat{{\beta}}$ by plugging in the estimates $\hat{\sigma}_{A}^{2}$ and $\hat{\sigma}_{B}^{2}$,
+$\hat{\mathrm{var}}(\hat{{\beta}})=(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})\,\hat{\mathrm{var}}(\hat{{\gamma}})=(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})\,{J}_{\text{all}}^{-1}(\hat{{\gamma}})\hat{V}_{\text{all}}(\hat{{\gamma}}){J}_{\text{all}}^{-1}(\hat{{\gamma}}).$ (13)
 A limitation of this approach is that it neglects the uncertainty in the estimation of the variance components: Although we do not expect a substantial impact in high-dimensional applications, it could be possible to adjust the (13) for the variability of the variance components through bootstrapping the row-wise and column-wise estimates $\hat{\tau}^{2}_{A}$ and $\hat{\tau}^{2}_{B}$.
 3 Simulations
-In this section, we simulate from the probit model with crossed random effects (1) and compare the performance of the ARC estimator with the traditional estimator obtained by maximizing the first-order Laplace approximation of the likelihood.
-A further method considered is given by an infeasible oracle estimator that uses the unknown true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to estimate the regression parameters as $\hat{\bm{\beta}}_{\mathrm{oracle}}=\hat{\bm{\gamma}}(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ The ARC method instead corrects $\hat{\bm{\gamma}}$ using estimates of the variance components.
+In this section, we simulate from the probit model with crossed random effects (1) and compare the performance of the all-row-column estimator with the traditional estimator obtained by maximizing the first-order Laplace approximation of the likelihood.
+A further method considered is an ‘infeasible oracle estimator’ that uses the unknown true values of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to estimate the regression parameters as $\hat{{\beta}}_{\mathrm{oracle}}=\hat{{\gamma}}(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ The all-row-column method instead corrects $\hat{{\gamma}}$ using estimates of the variance components.
 All methods were implemented in the R language.
 The package TMB (Kristensen et al., 2016) was used for the Laplace approximation, with the nlminb optimization function employed for its maximization.
-We did not use the popular glmer function from the R package lme4 (Bates et al., 2015) because TMB is substantially more computationally efficient and thus allowed us to compare our method to the first-order Laplace approximation at larger dimensions than it would otherwise be possible with glmer.
+We did not use the popular glmer function from the R package lme4 (Bates et al., 2015) because the current version of TMB is substantially more computationally efficient and thus allowed us to compare our method to the first-order Laplace approximation at larger dimensions than it would otherwise be possible with glmer.
 The row- and column-wise likelihoods were coded in C++ and integrated in R with Rcpp (Eddelbuettel, 2013), and optimized by Brent’s method as implemented in the optimise function.
-The code is available in the public repository https://github.com/rugbel/arcProbit.
 3.1 Simulation settings
@@
 The third factor is whether the random effect variances are set at a high level or at a low level.
-Given $R$ and $C$, the set $\mathcal{S}$ is obtained by IID Bernoulli sampling with probability $N/(RC)$.
-This makes the attained value of $N$ random but with a very small coefficient of variation.
-We denote by $R=N^{\rho}$ and $C=N^{\kappa}$ the number of rows and columns in the data, as a power of the total sample size $N$. The two levels of the balance factor are
- Balanced
-$\rho=\kappa=0.56$, and
- Imbalanced
-$\rho=0.88$ and $\kappa=0.53$.
-This balanced case was already considered in Ghosh et al.
-(2022b).
+Given $R$ and $C$, the set $\mathcal{S}$ is obtained by independent and identically distributed Bernoulli sampling with probability $1.27\times N/(RC)$.
+The value $1.27$ is the largest one for which Ghosh et al.
+(2022a) could prove that backfitting takes $O(1)$ iterations.
+This sampling makes the attained value of $N$ random but with a very small coefficient of variation.
+We denote by $R=N^{\rho}$ and $C=N^{\kappa}$ the number of rows and columns in the data, as a power of the total sample size $N$.
+The two levels of the balance factor are thereafter termed ‘balanced’ with $\rho=\kappa=0.56$ and ‘imbalanced’ with $\rho=0.88$ and $\kappa=0.53$.
+The balanced case was used in Ghosh et al.
+(2022b), and the imbalanced case is similar to the Stitch Fix data in §4.
 Because $\rho+\kappa>1$, the fraction of possible observations in the data is $N/(RC)=N^{1-\rho-\kappa}\to 0$ as $N\to\infty$ providing asymptotic sparsity in both cases.
 While the first choice has $R/C$ constant, the second choice has $R/C\to\infty$ with $N$.
-We believe that this asymptote is a better description of our motivating problems than either a setting with $C$ fixed as $R\to\infty$ or the setting common in random matrix theory (Edelman and Rao, 2005) where $R$ and $C$ diverge with $R/C$ approaching a constant value.
+We believe that this asymptote is a better description of our motivating problems than either a setting with $C$ fixed as $R\to\infty$ or the setting common in random matrix theory (Edelman & Rao, 2005) where $R$ and $C$ diverge with $R/C$ approaching a constant value.
+Notice that when $\rho=0.88$ and $\kappa=0.53$ the condition $\max(\rho+2\kappa,2\rho+\kappa)<2$ invoked for the estimator in Ghosh et al.
+(2022b) does not hold.
 We considered seven predictors generated from a multivariate zero-mean normal distribution with a covariance matrix $\Sigma$ corresponding to an autocorrelation process of order one, so that the entry $(k,l)$ of $\Sigma$ is $\phi^{|k-l|}$.
 We set $\phi=0.5$ in all the simulations.
-We always used the intercept $\beta_{0}=-1.2$ because in our applications $\Pr(Y=1)<1/2$ is typical.
-For the predictor coefficients we considered the following two choices
- Null
-$\beta_{\ell}=0$ for $\ell=1,\ldots 7$, and
- Linear
-$\beta_{\ell}=-1.2+0.3\ell$ for $\ell=1,\dots,7.$
-The first setting is a null one where $\bm{x}$ is not predictive at all, while the second setting has modestly important nonzero predictors whose values are in linear progression.
-The two choices for the variance component parameters are
- High variance
-$\sigma_{A}=1$ and $\sigma_{B}=1$, and
- Low variance
-$\sigma_{A}=0.5$ and $\sigma_{B}=0.2$.
+We always used the intercept $\beta_{0}=-1.2$ because in our applications $\mathrm{pr}(Y=1)<1/2$ is typical.
+For the predictor coefficients we considered two choices termed ‘null’ with $\beta_{\ell}=0$ and ‘linear’ with $\beta_{\ell}=-1.2+0.3\ell$ $(\ell=1,\dots,7).$ The first setting is a null one where ${x}$ is not predictive at all, while the second setting has modestly important nonzero predictors whose values are in linear progression.
+The two choices for the variance component parameters are termed ‘high variance’ with $\sigma_{A}=1$ and $\sigma_{B}=1$ and ‘low variance’ with $\sigma_{A}=0.5$ and $\sigma_{B}=0.2$.
 We chose the first setting to include variances higher than what is typically observed in applications.
-The second setting is closer to what we have seen in data such as that in Section 5.
+The second setting is closer to what we have seen in data such as that in §4.
 We represent the eight settings with mnemonics as shown in Table 1.
@@
 For each of these 13 sample sizes and each of the eight settings, we simulated 1000 data sets.
+As suggested by a referee, we experimented with different values for the number of quadrature nodes $k$ to approximate the univariate integrals of the row- and column-likelihoods; see also the recent work Bilodeau et al.
+(2024).
+The value $k=1,$ which corresponds to the Laplace approximation, produced estimates of $\sigma_{A}$ and $\sigma_{B}$ affected by substantial downward bias even at large $N$.
+The bias disappeared by increasing $k$.
+With $k=5,$ we obtained results not affected by bias and essentially indistinguishable from those with $k=25$.
+Therefore, the results discussed in the rest of this section were calculated with $k=5$ nodes.
 Graphs comparing the computational costs, the statistical properties and the scalability of the three estimation methods for all eight settings are reported in the Supplementary Materials.
@@
 3.2 Computational cost
-Computational cost is the first and most critical quantity to consider since methods that are too expensive cannot be used, regardless of their quality.
 Figure 2 shows the average computation times for the three methods, obtained on a 16-core 3.5 GHz AMD processor equipped with 128 GB of RAM.
-It also shows regression lines of log cost versus $\log(N)$, marked with the regression slopes.
-The ARC and oracle method’s slope are both very nearly $1$ as expected.
+It also shows regression lines of log cost versus $\log N$, marked with the regression slopes.
+The all-row-column and oracle method’s slope are both very nearly $1$ as expected.
 The Laplace method slope is clearly larger than $1$ and, as noted above, we curtailed the sample sizes used for that case to keep costs reasonable.
-Similar estimated computational costs were obtained for the other seven settings as shown in the Supplementary Material are summarized in Table 2 which shows estimated computational cost rates for all eight settings.
-The ARC slope is consistently very close to $1$.
+If we extrapolate the Laplace cost to $N=10^{8}$, comparable to the Netflix data (Bennett & Lanning, 2007), then the cost grows past 12.9 days while the all-row-column cost grows only to about 45 minutes.
+The computational cost of all-row-column can be further reduced with parallel computing by distributing the approximations of the R+C one-dimensional integrals across multiple cores.
+Figure 2: The left panel shows time in seconds versus sample size $N$ for the Imb-Nul-Hi setting for all three methods: Laplace, all-row-column and the Oracle.
+The right panel is the same plot with times displayed on the $\log_{10}$ scale.
+Similar estimated computational costs were obtained for the other seven settings as summarized in Table 2 which shows estimated computational cost rates for all eight very close to $1$.
 The oracle slope is consistently close to $1$ in imbalanced settings and somewhat bigger than $1$ in balanced ones.
@@
 It tends to be higher for balanced settings although one of the imbalanced cases also has a large slope.
+Setting Oracle slope ARC slope Laplace slope
+Bal-Nul-Hi 1.06 1.01 1.30
+Bal-Nul-Lo 1.05 1.01 1.24
+Bal-Lin-Hi 1.06 1.00 1.29
+Bal-Lin-Lo 1.07 0.99 1.26
+Imb-Nul-Hi 0.98 1.00 1.23
+Imb-Nul-Lo 0.98 1.00 1.13
+Imb-Lin-Hi 1.02 0.99 1.15
+Imb-Lin-Lo 0.98 1.00 1.15
+Table 2: Computational cost for all eight settings. Linear regression slopes for $\log(\text{CPU seconds})$ versus $\log(N).$
 We have investigated the data behind the oracle slopes, but cannot yet explain the mild superlinearity we sometimes see.
@@
 The anomaly is concentrated in the balanced simulations, but knowing this has not been enough to identify the cause.
-Setting Oracle slope ARC slope Laplace slope
-Bal-Nul-Hi 1.06 1.01 1.30
-Bal-Nul-Lo 1.05 1.01 1.24
-Bal-Lin-Hi 1.06 0.99 1.29
-Bal-Lin-Lo 1.07 0.99 1.26
-Imb-Nul-Hi 0.98 0.99 1.23
-Imb-Nul-Lo 0.98 1.00 1.13
-Imb-Lin-Hi 1.02 0.99 1.15
-Imb-Lin-Lo 0.98 0.99 1.15
-Table 2: For all eight settings, this table shows the observed computational cost as a rate in $N$, fit by linear regression of the logarithm of average CPU seconds on $\log(N)$.
-Figure 1: Time in seconds versus sample size N for the Imb-Nul-Hi setting for all three algorithms: Laplace, ARC and the Oracle.
 3.3 Regression coefficient estimation
@@
 The intercept poses a challenge because it is somewhat confounded with the random effects.
-For instance, if we replace $\beta_{0}$ by $\beta_{0}+\lambda$ while replacing $a_{i}$ by $a_{i}-\lambda$ then $Y_{ij}$ are unchanged.
-Large $\lambda$ would change $\bar{a}=(1/R)\sum_{i=1}^{R}a_{i}$ by an implausible amount that should be statistically detectable given that $a_{i}\sim\mathcal{N}(0,\sigma^{2}_{A})$.
-On the other hand $|\lambda|=O(\sigma_{A}/\sqrt{R})$ would be hard to detect statistically.
+For instance, if we replace $\beta_{0}$ by $\beta_{0}+\lambda$ while replacing $a_{i}$ by $a_{i}-\lambda$ then the $Y_{ij}$ are unchanged.
+Large $\lambda$ would change $\bar{a}=(1/R)\sum_{i=1}^{R}a_{i}$ by an implausible amount that should be statistically detectable given that $a_{i}\sim{N}(0,\sigma^{2}_{A})$.
+On the other hand $|\lambda|=O(\sigma_{A}R^{-1/2})$ would be hard to detect statistically.
 The other regression parameters are not similarly confounded with main effects in our settings.
-We note that when a predictor is a categorical variable that is either a function of the row index $i$ or of the column index $j$, then it can be subject to some confounding similar to the intercept’s confounding.
-This was observed in Ghosh et al.
-(2022b).
-Because of the confounding described above, we anticipate that the true MSE rate for the intercept cannot be better than $\mathcal{O}(\min\{R,C\}))$ which is $\mathcal{O}(N^{-0.53})$ in our imbalanced settings and $\mathcal{O}(N^{-0.56})$ in our balanced settings.
-For the other coefficients $\mathcal{O}(N^{-1})$ is not ruled out by this argument.
-Figures 2 and 3 reports the mean square errors for the intercept and the coefficient of the first predictor estimated for different sample sizes, for the three estimators under study in the Imb-Nul-Hi setting.
-We report here the plot only for the first predictor because the mean squared errors of the estimates of the seven regression coefficients were essentially equivalent.
-All three estimators show an MSE very close to $\mathcal{O}(N^{-1})$ for $\beta_{1}$.
-Where we anticipated an MSE no better than $\mathcal{O}(N^{-0.53})$ (imbalanced) or $\mathcal{O}(N^{-0.56})$ (balanced) for the intercept we saw slightly better MSEs with slopes between $-0.57$ and $-0.60$, confirming our expectation that the intercept would be harder to estimate than the regression coefficients.
-Figure 2: Mean squared errors for coefficient estimates in the Imb-Nul-Hi setting.
-The top row is for $\beta_{0}$ and the bottom row is for $\beta_{1}$.
-Each method is shown with a line and solid points and fitted reference lines for the other two methods (that overlap somewhat).
+Ghosh et al.
+(2022b) saw that a categorical predictor that is a function of just the row index $i$ or just the column index $j$ brings a similar confounding.
+Because of the confounding described above, we anticipate that the true mean square error rate for the intercept cannot be better than $O\{\min(R,C)\}$ which is $O(N^{-0.53})$ in our imbalanced settings and $O(N^{-0.56})$ in our balanced settings.
+For the other coefficients $O(N^{-1})$ is not ruled out by this argument.
+Supplementary Material Figure S5 reports the mean square errors for the intercept and the coefficient of the first predictor estimated for different sample sizes, for the three estimators under study in the Imb-Nul-Hi setting.
+We report there the plot only for the first predictor because the mean square errors of the estimates of the seven regression coefficients were essentially equivalent.
+All three estimators show a mean square error very close to $O(N^{-1})$ for $\beta_{1}$.
+Where we anticipated a mean square error no better than $O(N^{-0.53})$ (imbalanced) or $O(N^{-0.56})$ (balanced) for the intercept we saw slightly better mean square errors with slopes between $-0.57$ and $-0.61$, confirming our expectation that the intercept would be harder to estimate than the regression coefficients.
 3.4 Variance component estimation
 Here we present the estimation errors in the variance component parameters $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$.
-The oracle method is given the true values of these parameters and so the comparison is only between ARC and the Laplace method.
+The oracle method is given the true values of these parameters and so the comparison is only between all-row-column and the Laplace method.
 For the variance parameter $\sigma^{2}_{A}$, the data only have $R$ levels $a_{1},\dots,a_{R}$.
-If they were observed directly, then we could estimate $\sigma^{2}_{A}$ by $(1/R)\sum_{i=1}^{R}a_{i}^{2}$ and have an MSE of $\mathcal{O}(R^{-1})$.
+If they were observed directly, then we could estimate $\sigma^{2}_{A}$ by $(1/R)\sum_{i=1}^{R}a_{i}^{2}$ and have a mean square error of $O(R^{-1})$.
 In practice, the $a_{i}$ are obscured by the presence of the signal, the noise $\varepsilon_{ij}$ and the other random effects $b_{j}$.
-Accordingly, the best rate we could expect for $\sigma^{2}_{A}$ is $\mathcal{O}(R^{-1})=\mathcal{O}(N^{-\rho})$ and the best we could expect for $\sigma^{2}_{B}$ is $\mathcal{O}(C^{-1})=\mathcal{O}(N^{-\kappa})$.
-Figure 3 shows the MSE for estimation of $\sigma_{A}$ and $\sigma_{B}$ for ARC and Laplace methods in the Imb-Nul-Hi setting.
-Due to the imbalance, our anticipated rates are $\mathcal{O}(N^{-0.88})$ for $\sigma_{A}$ and $\mathcal{O}(N^{-0.53})$ for $\sigma_{B}$.
-The ARC method does slightly better than these rates.
+Accordingly, the best rate we could expect for $\sigma^{2}_{A}$ is $O(R^{-1})=O(N^{-\rho})$ and the best we could expect for $\sigma^{2}_{B}$ is $O(C^{-1})=O(N^{-\kappa})$.
+Figure 3 shows the mean square error for estimation of $\sigma_{A}$ and $\sigma_{B}$ for all-row-column and Laplace methods in the Imb-Nul-Hi setting.
+Due to the imbalance, our anticipated rates are $O(N^{-0.88})$ for $\sigma_{A}$ and $O(N^{-0.53})$ for $\sigma_{B}$.
+The all-row-column method does slightly better than these rates.
 The Laplace method attains nearly this predicted rate for $\sigma_{B}$ but does much worse for $\sigma_{A}$.
 We can understand both of these discrepancies in terms of biases, described next.
-Figure 4 shows boxplots for the parameter estimates of $\sigma_{A}$ and $\sigma_{B}$ with the ARC and Laplace methods.
-We can compare the center of those boxplots to the reference line at the true parameter values and see that for the ARC estimates there is a bias decreasing at a faster rate than the width of the boxes.
-This explains the slightly better than predicted rates that we see for ARC.
+Figure 3: Mean square errors and estimated convergence rates for $\sigma_{A}$ and $\sigma_{B}$ for all-row-column and Laplace methods in setting Imb-Nul-Hi.
+The dotted line in the plot for the Laplace method is the convergence rate of all-row-column.
+In each plot there is a reference
+Figure 4 shows boxplots for the parameter estimates of $\sigma_{A}$ and $\sigma_{B}$ with the all-row-column and Laplace methods.
+We can compare the center of those boxplots to the reference line at the true parameter values and see that for the all-row-column estimates there is a bias decreasing at a faster rate than the width of the boxes.
+This explains the slightly better than predicted rates that we see for all-row-column.
 Instead the Laplace method has a substantial bias that only decreases very slowly as $N$ increases, giving Laplace a worse than expected rate, especially for $\sigma_{A}$.
-In order to understand the difficulties of the Laplace approximation method we turn to the literature on Laplace approximations where the dimension of the integral grows with the sample size.
-The crossed random effects integral has dimension $R+C$ for $N$ observations.
-The largest number of parameters in any of our simulations was for the imbalanced setting with $R+C=\mathcal{O}(N^{0.88})$ and this is where the Laplace approximation had its worst performance.
-Section 6 of Shun and McCullagh (1995) suggests that a sufficient condition for the Laplace approximation to be reliable is, in our notation, that $R+C=o(N^{1/3})$.
-This is not possible in a sparse setting with $RC\gg N$.
-Ogden (2021) notes that Shun and McCullagh (1995) use an assumption on the growth rate of the derivatives of the exponent in the Laplacian integrand and says that assumption is not realistic for generalized linear mixed models.
-She then gives an analysis of the hierarchical setting which we believe is much simpler than the present crossed setting.
-For a hierarchical model with our row random effects but no column random effects, the likelihood is a product of $R$ one dimensional integrals.
-The logarithm of the $i$’th integrand is a sum of $N_{i\text{\tiny$\bullet$}}$ terms.
-She notes that the error in a first order Laplace approximation is $\mathcal{O}(\sum_{i}N_{i\text{\tiny$\bullet$}}^{-1})$.
-The smallest that could be with $N$ observations in $R$ columns is $R(N/R)^{-1}=R^{2}/N$, a quantity that is $\mathcal{O}(N^{2\rho-1})=\mathcal{O}(N^{0.76})$ in our unbalanced setting.
-Things improve with higher order approximations.
-For the definition of $k$’th order in Ogden (2021), equation (10) there gives an error bound of $\mathcal{O}(\sum_{i}N_{i\text{\tiny$\bullet$}}^{-k})$.
-Ogden (2021) remarks that this bound might be conservative.
-All of our simulations had $\sum_{i}N_{i\text{\tiny$\bullet$}}^{-1}$ increasing with $N$ but only in the imbalanced null cases did it look like the Laplace method might be inconsistent (due to bias).
-For the approximation of one-dimensional integrals in ARC an accurate numerical approximation was employed and we saw empirical convergence rates close to those we anticipated.
-Figure 3: MSEs and estimated convergence rates for $\sigma_{A}$ and $\sigma_{B}$ for ARC and Laplace methods in setting Imb-Nul-Hi.
-In each plot there is a reference line for the covergence rate of the other method.
-Figure 4: Boxplots of $\sigma_{A}$ and $\sigma_{B}$ estimates for the ARC and Laplace methods in the setting Imb-Nul-Hi.
-There are horizontal reference lines at the true parameter values.
-The ARC estimates have negative bias at small $N$ while the Laplace estimates have a strong extreme negative bias for all $N$.
+Figure 4: Boxplots of $\sigma_{A}$ and $\sigma_{B}$ estimates for the all-row-column and Laplace methods in the setting Imb-Nul-Hi. There are horizontal reference lines at the true parameter values.
 3.5 Other settings
-The simulation results for all eight settings, are reported in full in the Supplementary Figures S5–S44.
+The simulation results for all eight settings, are reported in full in the Supplementary Figures S1–S24.
 We already discussed computational cost in eight settings based on Table 2.
 Here we make brief accuracy comparisons.
-We compare our proposed ARC method to the oracle method which is infeasible because it requires knowledge of $\sigma_{A}$ and $\sigma_{B}$ and to the Laplace method which becomes infeasible for large $N$ because it does not scale as $\mathcal{O}(N)$.
-In most settings and for most parameters, the oracle method was slightly more accurate than ARC.
+We compare our proposed all-row-column method to the oracle method which is infeasible because it requires knowledge of $\sigma_{A}$ and $\sigma_{B}$ and to the Laplace method which becomes infeasible for large $N$ because it does not scale as $O(N)$.
+In most settings and for most parameters, the oracle method was slightly more accurate than all-row-column.
 That did not always hold.
-For instance in setting Bal-Lin-Hi (Figure S14) we see that ARC was slightly more accurate than the oracle method for $\beta_{1}$.
-The figures in the Supplementary material show some cases where ARC has a slight advantage over the Laplace method and some where it has a slight disadvantage.
+We see that all-row-column was slightly more accurate than the oracle method for $\beta_{1}$ in three of the eight settings, namely Bal-Lin-Hi (Figure S3), Bal-Lin-Lo (Figure S4) and Imb-Lin-Hi (Figure S7).
+The figures in the Supplementary material show some cases where all-row-column has a slight advantage over the Laplace method and some where it has a slight disadvantage.
 There are a small number of cases where the Laplace method has outliers at $N=10^{3}$ that cause the linear regression slope to be questionable.
-These are present in setting Imb-Lin-Hi (Figure S20).
-However the attained MSEs at $N=10^{5}$ do not differ much between ARC and Laplace in that setting.
-Our conclusion is that compared to the Laplace method, ARC is scalable and robust.
-The Supplementary Material Section S2 makes a comparison of ARC with the maximum pairwise likelihood estimator of Bellio and Varin (2005).
+These are present in the estimates of $\sigma_{A}$ and $\sigma_{B}$ for setting Imb-Lin-Hi (Figure S23).
+However the attained mean square errors at $N=10^{5}$ do not differ much between all-row-column and Laplace in that setting (Figure S15).
+Our conclusion is that compared to the Laplace method, all-row-column is scalable and robust.
+The Supplementary Material §S3 makes a comparison of all-row-column with the maximum pairwise likelihood estimator of Bellio & Varin (2005).
+Their pairwise likelihood involves all the pairs of correlated observations, that is those pairs that share the row- or the column-random effect.
+The pairwise likelihood is not computationally attractive.
+The number of pairs of data points far exceeds the $O(N)$ constraint.
+In a setting with $R$ rows, the average row has $N/R$ elements in it.
+If all the rows had that many elements then the number of pairs would be $\Omega\{R(N/R)^{2}\}=\Omega(N^{2}/R)$.
+Unequal numbers of observations per row can only increase this count.
+As a result the cost must be $\Omega\{\max(N^{2}/R,N^{2}/C)\}=\Omega\{\max(N^{2-\rho},N^{2-\kappa})\}$, so the cost cannot be $O(N)$.
 The comparison is made for the Imb-Nul-Hi setting that we have been focusing on.
-From Section 2, the cost of pairwise composite likelihood is $\Omega(\max(N^{2}/R,N^{2}/C))=\Omega(\max(N^{2-\rho},N^{2-\kappa}))$.
-For the imbalanced setting this is $\Omega(N^{2-0.53})=\Omega(N^{1.47}).$ We see in Section S2 that the empirical cost of the pairwise composite likelihood grows as $N^{1.42}$, close to the predicted rate.
-The pairwise method attains very similar parameter estimation accuracy to ARC.
-The most important difference for this example is that ARC costs $\mathcal{O}(N)$ while pairwise likelihood is far more expensive and does not scale to large data sets.
+For that imbalanced setting, the cost of the pairwise likelihood is $\Omega(N^{2-0.53})=\Omega(N^{1.47}).$ We see in §S3 that the empirical cost of the pairwise composite likelihood grows as $N^{1.46}$, close to the predicted rate.
+The pairwise method attains very similar parameter estimation accuracy to all-row-column.
+The most important difference for this example is that all-row-column costs $O(N)$ while pairwise likelihood is far more expensive and does not scale to large data sets.
 4 Application to the Stitch Fix Data
-In this section we illustrate our ARC method on a data set from Stitch Fix. Quoting from Ghosh et al. (2022b):
-Stitch Fix is an online personal styling service.
+In this section we illustrate our all-row-column method on a data set from Stitch Fix. As described in Ghosh et al. (2022b):
+“Stitch Fix is an online personal styling service.
 One of their business models involves sending customers a sample of clothing items.
@@
 It is also from 2015.
-While it does not describe their current business, it is a valuable data set for illustrative purposes.
+While it does not describe their current business, it is a valuable data set for illustrative purposes.”
 The Stitch Fix data consists of $N=5{,}000{,}000$ ratings from $R=744{,}482$ clients on $C=3{,}547$ items.
@@
 The predictor variables we used are listed in Table 3.
+There is one block of client predictors followed by a block of item predictors.
 Some of the categorical variable levels in the data had only a small number of levels.
@@
 vegetable fiber
-Item fit item fit fitted
+Item fit fit of the clothing item fitted
 loose or oversized
 straight or tight
-Table 3: Predictors available in the Stitch Fix data. There is one block of client predictors followed by a block of item predictors.
-As anticipated in Section 2, we assume missing data at random and have removed rows with any missing entries.
-That leaves us with $N=4{,}965{,}960$ of the original $5{,}000{,}000$ ratings.
-We have data from $741{,}221$ clients on $3{,}523$ items.
+Table 3: Predictors available in the Stitch Fix data.
+Some of the observations with $Z_{ij}=1$ were nonetheless incomplete with a few missing entries.
+Deleting them left us with $N=4{,}965{,}960$ ratings from $741{,}221$ clients on $3{,}523$ items.
 The data are not dominated by a single row or column.
@@
 In a business setting one would fit and compare a wide variety of different binary regression models in order to understand the data.
-Our purpose here is to study large scale probit models with including crossed random effects and so we choose just one model for illustration, possibly the first model of many that one would consider.
+Our purpose here is to study large-scale probit models with including crossed random effects and so we choose just one model for illustration, possibly the first model of many that one would consider.
 We consider the probit model with crossed random effects whose fixed effects are specified according to the symbolic model formula:
@@
 The first level of each categorical predictor in alphabetical order (Table 3) is used as the reference level in fitting the model.
-Table 4 reports 1) the maximum likelihood estimates of the regression parameters under a marginal probit model that ignores the customer and item heterogeneity, and 2) the ARC estimates for the probit model with two crossed random effects for the customers and items.
-The z values reported in the table are computed with the observed information for the marginal probit model and with the two-way cluster-robust sandwich estimator described in Section 3.2.
-The ARC estimates of the variance components are $\hat{\sigma}_{A}=0.53$ and $\hat{\sigma}_{B}=0.34$.
-As expected, ignoring the customer and item heterogeneity lead to a large underestimation of the uncertainty in the parameter estimates and thus in the marginal probit all the predictors are strongly significant given the very large sample size.
-Conversely, the cross-random effects model takes into account the sources of heterogeneity and leads us to learn that the item fit is not a significant predictor of the top rank and that items made by vegetable fibers are less likely to be ranked top compared to clothes made by artificial fibers.
-While we have emphasized the naive variance estimation of the marginal probit model, Table 4 also shows the attenuation bias.
-The naive model systematically underestimates the absolute value of the coefficients while simultaneously overestimating their statistical significance.
-Marginal probit Random-effects probit
+Table 4 reports 1) the maximum likelihood estimates of the regression parameters under a naïve probit model that ignores the customer and item heterogeneity, and 2) the all-row-column estimates for the probit model with two crossed random effects for the customers and items.
+The random effects probit parameter estimate $\hat{\beta}$ equals the naïve probit estimate $\hat{\gamma}$ multiplied by $(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2}$.
+Using adaptive Gauss-Hermite quadrature we obtained estimates $\hat{\sigma}_{A}\doteq 0.53$ and $\hat{\sigma}_{B}\doteq 0.34$.
+Following what we learned in the simulation studies, the estimates on all rows and columns were calculated with $k=5$ quadrature nodes.
+An earlier conservative calculation used $k=28$ but $k=5$ gave indistinguishable results.
+The z values reported in the table are computed with the observed information for the naïve probit model and with the two-way cluster-robust sandwich estimator described in §2.3.
+As expected, ignoring the customer and item heterogeneity leads to a large underestimation of the uncertainty in the parameter estimates and thus in the naïve probit all the predictors are strongly significant given the very large sample size.
+Conversely, the crossed random effects model takes into account the sources of heterogeneity and leads us to learn that the item fit is not a significant predictor of the top rank and that items made by vegetable fibers are less likely to be ranked top compared to clothes made by artificial fibers.
+Naïve probit Random-effects probit
 Variable ${}^{\text{\textdagger}}$ z value p value ${}^{\text{\textdagger}}$ z value p value
@@
 […]
-Material leather or animal --12.9 --12.99 $<0.001$-15.2 --1.57 0.116
+Material leather or animal --12.9 --12.99 $<0.001$-15.2 --1.57 0.1160.116{0.116}0.116
 […]
@@
 ${}^{\text{\textdagger}}$ll the estimated predictor parameters are multiplied by 100.
-Nonsignificant p values are marked in bold.
 See Table 3 for a description of the predictors.
 The first level of each categorial predictor in alphabetic order is used as reference level.
-Figure 5 compares the two-way cluster robust sandwich standard errors with 1) the naive standard errors from a probit fit ignoring dependence between items and customers and 2) the pigeonhole nonparametric bootstrap standard errors of Owen (2007) that do not assume correct model specification, as mentioned in Section 3.2.
-The figure shows how closely the sandwich and pigeonhole standard errors agree, while the naive standard errors are much smaller.
-The naive values correspond to variances underestimated by factors ranging from 4 to just over 1240, depending on the parameter and only slightly on whether we use sandwich or pigeonhole estimates of the coefficient variances.
+Supplementary Material Figure S29 compares the two-way cluster robust sandwich standard errors with 1) the standard errors from the naïve probit fit multiplied by $(1+\hat{\sigma}^{2}_{A}+\hat{\sigma}^{2}_{B})^{1/2}\doteq 1.18$ to report them in the fixed effects scale of the probit model with crossed random effects and 2) the pigeonhole nonparametric bootstrap standard errors of Owen (2007) that do not assume correct model specification, as mentioned in §2.3.
+The Figure S29 shows how closely the sandwich and pigeonhole standard errors agree.
+The naïve standard errors, which ignore dependence between items and customers, correspond to variances underestimated by factors ranging from 3 to 954, depending on the parameter and only slightly on whether we use sandwich or pigeonhole estimates of the coefficient variances.
+Thus ignoring the dependencies from correlated data makes an enormous difference here.
 These standard errors are reported in Table S1 of the Supplementary Material.
 We also computed parametric bootstrap standard errors (not shown).
-Those were in between the naive ones and the nonparametric ones.
-Figure 5: Stitch Fix Application.
-Comparison between the two-way cluster robust sandwich standard errors (x axis), the naive GLM standard errors (y axis, triangles) and the Pigeonhole nonparametric bootstrap standard errors (y axis, circles).
-The gray reference line is the first quadrant bisector $y=x$.
-All standard errors are multiplied by 100.
-6 Bibliographic Notes
-In this section we review some related literature on the difficulties that crossed random effects bring.
-McCullagh (2000) points out a key difference between the crossed and nested random effects.
-Even in an intercept-only setting with $Y_{ij}=\mu+a_{i}+b_{j}+\varepsilon_{ij}$ and all $R\times C$ values observed, there is no correct bootstrap for the variance of the simple average of the $Y_{ij}$.
-Resampling rows and columns independently is mildly conservative in this case.
-That mild conservatism also extends to some sparse settings (Owen, 2007).
-Gao and Owen (2017) provide an $\mathcal{O}(N)$ algorithm to consistently estimate an intercept plus variance components.
-Gao and Owen (2020) provides an $\mathcal{O}(N)$ algorithm for linear regression with crossed random effects; that algorithm is less efficient than generalized least squares but the variance estimates from it properly account for all of the random effects.
-They also show that computing the Gaussian likelihood for crossed random effects at just one parameter value costs $\Omega(N^{3/2})$.
-Ghosh et al.
-(2022a) provide a backfitting algorithm for generalized least squares that costs $\mathcal{O}(N)$ per iteration and they give asymptotic conditions under which the number of iterations is $\mathcal{O}(1)$.
-Ghandwani et al.
-(2023) extended the algorithm to a linear mixed model that include random slopes.
-Ghosh et al.
-(2022b) consider logistic regression with random intercepts adapting an integral approximation from Schall (1991).
-Their iterations cost $\mathcal{O}(N)$ and empirically the algorithm runs in $\mathcal{O}(N)$ time, but there is no proven bound on the number of iterations required.
-The computational challenge from crossed random effects is also present for Bayesian methods.
-A plain Gibbs sampler costs $\mathcal{O}(N)$ per iteration but takes $\Omega(N^{1/2})$ iterations for the model in Gao and Owen (2017).
-Papaspiliopoulos et al.
-(2023) derived a scalable MCMC algorithm for generalized linear models with crossed-random effects that extends previous work on a collapsed Gibbs sampler described in Papaspiliopoulos et al.
+Those were somewhat lower than the nonparametric standard errors as expected.
+In an application like that of Stitch Fix data, the typical goal is to make inference about the probability of ranking an item top for a specific customer and a specific item.
+Such evaluations also require estimating the customer ($a_{i}$) and item ($b_{j}$) random effects.
+The estimates of those random effects are a byproduct of the adaptive Gauss-Hermite quadrature used to approximate the row- and column-wise likelihoods.
+Figure S30 in the Supplement shows the distribution of the estimated customer and item random effects.
+5 Discussion
+In Theorem 2 we proved that there exists a root of the row (or column) likelihood equation that is a consistent estimator of $\tau^{2}_{A}$ (or $\tau^{2}_{B}).$ This form of consistency is commonly called Cramer consistency.
+It is the same notion of consistency that Jiang (2013) established for the maximum likelihood estimate.
+If one does not find Cramer consistency sufficient, then it is possible to construct estimators $\hat{\tau}^{2}_{A}$ and $\hat{\tau}^{2}_{B}$ that converge in probability to $\tau^{2}_{A}$ and $\tau^{2}_{B}$ as $N\to\infty$.
+A consistent estimator of $\tau^{2}_{A}$ can be obtained from one or more rows $i$ for which $N_{i\text{\tiny$\bullet$}}\to\infty$ and a similar approach works for $\tau^{2}_{B}$.
+In §S1.3 of the Supplementary Material we show how to construct such a consistent estimator and we give conditions under which the number of large rows will diverge to infinity as $N\to\infty$.
+See Theorem 3 there.
+We prefer our all-row-column estimator to an approach using just large rows, because it would be awkward to have to decide in practice which rows to use, and because we believe that there is valuable information in the other, smaller, rows.
+We have used a probit model instead of a logistic one because a Gaussian latent variable is a very natural counterpart to the Gaussian random effects that are the default in random effects models.
+That connection simplified our modeling and computation.
+Gibbons & Hedeker (1997) remark that “As in the case of fixed effect models, selection of probit versus logistic response functions appears to have more to do with custom or practice within a particular discipline than differences in statistical properties.” An extension to logistic regression is outside the scope of this paper.
+We are confident that our approach of combining multiple misspecified models will extend to other settings with Gaussian random effects and latent variables.
+The code we use already handles ordinal regression.
+Extensions to more than two effects or multivariate responses may well work similarly, but are outside the scope of this article.
+We conclude with some additional references about recent work on inference for data with a crossed design.
+Goplerud et al.
+(2025) developed a variational approximation for scalable Bayesian estimation using an appropriate relaxing of the mean-field assumption to avoid underestimation of posterior uncertainty in high dimensions.
+Xu et al.
+(2023) combine variational approximations and composite likelihoods that consider row-column decomposition in a similar way to ours.
+This approach is particularly convenient for Poisson and gamma regression models because in this case analytical calculations allow the approximation of one-dimensional integrals that appear in the composite likelihood to be avoided.
+In the binary case that we considered in our article, the approach of Xu et al.
+(2023) requires numerical integration and its consistency has not yet been established.
+Hall et al.
+(2020) consider message passing algorithms for generalized linear mixed models and their §6 includes a crossed effects binary regression.
+However it has $R=10$ and $C=6$ in our notation with $3$ replicates at each $(i,j)$ pair, and it did not address scalability.
+Ruli et al.
+(2016) proposed an improved version of the Laplace approximation to overcome the potential failure of the usual Laplace approximation and also illustrates it in the case of generalized linear models with crossed random effects in their §3.5.
+However, this proposal is not scalable, as illustrated by the numerical results included in Ruli et al.
 (2016).
-This Bayesian simulation method is more general than our proposal because it is not restricted to a particular type of response.
-Our proposal instead derives a frequentist solution to the important problem of consistent and scalable estimation of a regression model for cross-classified binary responses.
-Differently from the MCMC method developed by Papaspiliopoulos et al.
-(2023), our approach does not require the use of simulation since it is only necessary to compute a large number of one-dimensional integrals which we approximate through an adaptive quadrature carefully implemented to avoid the risk of accumulating integration errors.
-Other recent work has considered variational approximations to build computational efficient fitting of generalized linear mixed models with crossed-random effects.
-Goplerud et al.
-(2023) developed a variational approximation for scalable Bayesian estimation using an appropriate relaxing of the mean-field assumption to avoid underestimation of posterior uncertainty in high-dimensions.
-Parallel work on frequentist inference for generalized linear models with crossed-random effects has been conducted by Xu et al.
-(2023) by combining variational approximations and composite likelihoods that consider row-column decomposition in a similar way to ours.
-This approach is particularly convenient for Poisson and gamma regression models because in this case analytical calculations allow the approximation of one-dimensional integrals that appear in the composite likelihood to be avoided.
-However, the consistency of the Xu et al.
-(2023) estimation method for binary data has not been established yet, and only a conjecture for the logistic case has been formulated by the authors.
-Further investigation of the inferential properties of their proposal for the case of binary data seems warranted for a thorough comparison of the two methods.
-Hall et al.
-(2020) consider message passing algorithms for generalized linear mixed models and their Section 6 includes a crossed effecs binary regression.
-However it has $R=10$ and $C=6$ in our notation with $3$ replicates at each $(i,j)$ pair, and it did not address scalability.
 Bartolucci et al.
@@
 This approach shares the philosophy of our proposal but differs from our work in terms of model (latent discrete Markov variables vs crossed random effects), fitting procedure (EM algorithm vs direct maximization), data structure (balanced vs unbalanced and sparse) and motivating application (genomics vs e-commerce).
-7 Conclusions
-We have developed a method for crossed random effects for binary regression.
-It is a new version of composite likelihood that, by exploiting the connection between the Gaussian latent variables and the Gaussian random effects in the probit model, makes for very convenient factors in the composite likelihood.
-In simulations, we find that the computational cost of our ARC proposal is $\mathcal{O}(N)$ and its accuracy is competitive with and even more robust than a computationally infeasible Laplace approximation that has a cost which grow superlinearly with $N$.
-The mean squared errors of our method are very nearly $\mathcal{O}(\min\{R,C\}^{-1})$ for the intercept and $\mathcal{O}(N^{-1})$ for the other regression parameters.
-They are nearly $\mathcal{O}(R^{-1})$ and $\mathcal{O}(C^{-1})$ for the row- and column-random effect variance parameters, respectively.
-In the application to the e-commerce Stitch Fix data, a naive probit model finds that the coefficients of a model with 12 regression coefficients are all very statistically significant.
-After properly accounting for correlations due to random effects, we find that only eight of them remain significant at the 5% level.
-The naive analysis might lead to one following false leads.
-It can even lead one to confidently attribute incorrect signs to some parameters, even though $\mathrm{sign}(\gamma_{j})=\mathrm{sign}(\beta_{j})$.
-This can happen when $\mathrm{sign}(\hat{\gamma}_{j})\neq\mathrm{sign}(\gamma_{j})$ and the estimate of $\mathrm{var}(\hat{\gamma}_{j})$ is unrealistically small.
-One surprise in this work was the difficulty of getting good approximations to even the one dimensional Gaussian integrals that arise.
-We are grateful for the insights in Ogden (2021) and Shun and McCullagh (1995).
-We surmise that the difficulties in multidimensional integration for random effects models on big data sets must be even more severe than the univariate ones we saw.
-We have used a probit model instead of a logistic one because a Gaussian latent variable is a very natural counterpart to the Gaussian random effects that are the default in random effects models.
-That connection simplified our modeling and computation.
-As noted in the introduction, logit and probit link functions are nearly proportional outside of tail regions.
-Gibbons and Hedeker (1997) remark that “As in the case of fixed effect models, selection of probit versus logistic response functions appears to have more to do with custom or practice within a particular discipline than differences in statistical properties.” Thus, while a logistic regression approach is outside the scope of the present paper we think that an extension to the logistic case is worth pursuing and that our present findings provide a good starting point.
-If we had used a logistic instead of the probit model, then the all-likelihood would have required us to approximate further $N$ univariate integrals with respect to a convolution of the Gaussian and logistic probability density functions.
-That would be much more expensive than the probit model costs, although still within the $\mathcal{O}(N)$ constraint.
-Another approach to extending the ARC method to logistic regression uses the fact that the logistic distribution is a scale mixture of the standard normal distribution (Stefanski, 1991).
-This property is often used in Bayesian computations, see, for example, Narisetty et al.
-(2018).
 Supplementary Material
-The supplementary material contains proofs of the theorems of Section 3, additional simulation results discussed in Section 4, further references to the literature and a summary table of the binary regression standard errors for the Stitch Fix data.
-R code for replicating our results is available in the public repository https://github.com/rugbel/arcProbit.
+The supplementary material contains proofs of the theorems of §2, additional simulation results discussed in §3, further references to the literature, as well as a table and two plots about the Stitch Fix application mentioned in §4.
+The supplement is at
+https://artowen.su.domains/reports/ARC-final-supplement.pdf.
+R code for replicating our results is available in the public repository
+https://github.com/rugbel/arcProbit.
 Acknowledgements
+We are grateful to Silvia Bianconcini, Blair Bilodeau, Nancy Reid, Alex Stringer, Yanbo Tang and Matt Wand for helpful discussion about generalized linear mixed models, Laplace approximation and adaptive Gauss-Hermite quadrature.
 We thank Bradley Klingenberg and Stitch Fix for making this data set available.